### v1

# Bin Yang1,2, Alexandru Paul Condurache1,2

Abstract
3D scene understanding is a critical yet challenging task in autonomous driving, primarily due to the irregularity and sparsity of LiDAR data, as well as the computational demands of processing large-scale point clouds. Recent methods leverage the range-view representation to improve processing efficiency. To mitigate the performance drop caused by information loss inherent to the ”many-to-one” problem, where multiple nearby 3D points are mapped to the same 2D grids and only the closest is retained, prior works tend to choose a higher azimuth resolution for range-view projection. However, this can bring the drawback of reducing the proportion of pixels that carry information and heavier computation within the network. We argue that it is not the optimal solution and show that, in contrast, decreasing the resolution is more advantageous in both efficiency and accuracy. In this work, we present a comprehensive re-design of the workflow for range-view-based LiDAR semantic segmentation. Our approach addresses data representation, augmentation, and post-processing methods for improvements. Through extensive experiments on two public datasets, we demonstrate that our pipeline significantly enhances the performance of various network architectures over their baselines, paving the way for more effective LiDAR-based perception in autonomous systems. The code will be released based on the acceptance.
1 Introduction
LiDAR is one of the most common sensors for perception in autonomous driving. Semantic segmentation on LiDAR point clouds is essential for getting useful and reliable information of the surrounding 3D environment. To solve this 3D scene understanding task, many prior works propose to integrate deep learning techniques because of its remarkable advancements in the past few years. The publication of various annotated datasets [11, 3, 30] in the domain of autonomous driving further promotes research in the field.
In general, those methods can be categorized based on LiDAR data representation into point-based [47, 26, 39], voxel-based [49, 14] and projection-based methods [46, 24, 5, 48]. Both point- and voxel-based approaches typically require substantial computational resources due to the need to process data through networks with numerous 3D convolutional layers, intensive feature pre-processing, and deep architectures involving multiple downsampling and upsampling operations. These requirements can result in slow inference speeds, limiting their suitability for real-time applications. In contrast, rasterizing point cloud into range-view images [7] is more advantageous in fast and scalable LiDAR perception, because it allows the use of 2D operators for efficient computation and facilitates the transfer of knowledge from camera images [1, 18].
[…]
Nevertheless, learning from range-view representations can suffer from the performance drop caused by the ”many-to-one” conflict of adjacent points. To resolve the problem, most approaches maximize the azimuth resolution to make the range image more informative [24, 7, 5, 48, 13, 18]. We argue that this is suboptimal for the following reasons: 1) High overhead: increasing resolution significantly raises computational demand, limiting real-time performance. 2) Inefficient computation: the sparse nature of LiDAR data leads to many empty grid cells. Those unoccupied pixels introduce noise in the data and consume resources needlessly. 3) Limited informative gains: due to distortions caused by ego-motion and sensor-internal errors, points mapped to the same azimuthal location do not perfectly align with the laser beams. This misalignment implies that prioritizing both azimuth and elevation resolutions, rather than increasing image width alone, would yield a higher projection rate, as seen in our analysis in Fig. 2.
Figure 2: Statistics on SemanticKITTI [2]: 3D validity (proportion of projected points) with different azimuth (W) and elevation (H) resolutions. Comparable increases are observable when doubling azimuth and elevation resolution ($\Delta V_{azi},\Delta V_{ele}$).
Building on the aforementioned insights, we introduce FLARES, a novel training schema for range-view semantic segmentation of LiDAR point clouds. Rather than focusing on designing a new network architecture, FLARES targets to address the common limitations outlined above, hence, our proposal is generalizable to any range-view-based approaches. In essence, FLARES divides the point cloud into multiple sub-clouds, with only one projected onto a low-resolution image during training. During inference, all sub-clouds are projected and stacked along the batch dimension for processing. This approach increases the projection rate in both azimuth and elevation, thereby enriching informativeness of the range-view representation at lower cost.
Downscaling resolution, however, can exacerbate class imbalance in the dataset [7], potentially leading to overfitting during training. Another problem is decreasing 2D occupancy due to splitting the point cloud. To tackle these two incidental issues, we extend the pipeline with two additional data augmentation steps. Furthermore, we explore improvements in post-processing methods, a topic that has received minimal attention in previous works.
In summary, our contributions are as follows: 1) We introduce FLARES, a newly designed training schema for faster and more accurate LiDAR semantic segmentation. 2) We integrate two data augmentation techniques tailored to the new schema to enhance the network performance. 3) We propose a novel interpolation-based post-processing approach to resolve the ”many-to-one” problem more effectively. 4) We generalize this schema and all components across various range-view semantic segmentation networks, demonstrating superior performance over baselines on two different benchmarks.
2 Related Works
Point- and Voxel-based methods Some recent works [26, 47, 38] use raw point cloud data as direct network input, eliminating the need for post-processing after prediction. However, these methods often face high computational complexity and memory usage. To address these issues, Hu et al. [17] introduced sub-sampling and feature aggregation techniques for large-scale point clouds to reduce computational costs. Despite these efforts, performance degradation remains significant. Other works [49, 19] use 3D voxel grids as input, achieving point-based accuracy with reduced computational costs by utilizing sparse 3D convolutions [6]. Nonetheless, voxelization and de-voxelization steps continue to be time- and memory-intensive. Range-view-based methods To address inefficiencies, some prior works [36, 24, 37] convert the large-scale point cloud to panoramic range image through spherical projection and leverage image segmentation techniques for LiDAR data. SalsaNext [7] uses a Unet-like network with dilated convolutions to broaden receptive fields for more accurate segmentation, while Lite-HDSeg [28] introduces an efficient framework using a lite version of harmonic convolutions. Additionally, FIDNet [48] and CENet  [5] interpolate and concatenate multi-scale features with a minimal decoder for semantic prediction. These methods share the benefit of lightweight network design, significantly improving efficiency and enabling real-time applications. Nevertheless, they generally underperform 3D methods due to the “many-to-one” issue, where multiple points project to the same pixel. To offset the performance drop caused by the problem, some other recent works propose to use Vision Transformer (ViT) [9, 41, 35]. RangeViT [1] deploys standard ViT backbone as encoder, followed by a light-weight decoder for refining the coarse patch-wise ViT representations, while RangeFormer [18] utilizes a pyramid-wise ViT-encoder to extract multi-scale features from range images. ViTs offer higher model capacities and excel at capturing long-range dependencies by modeling global interactions between different regions, enhancing segmentation performance over traditional CNNs [8]. However, the quadratic computational complexity of self-attention mechanisms in ViTs introduces challenges in achieving an optimal balance between efficiency and accuracy. Training schema Highlighting inefficiencies stemming from the use of high-resolution range images, some methods have adopted compact networks, significantly reducing network capacity [48, 7, 5]. Unfortunately, high-resolution range images still demand substantial memory, which restricts scalability in terms of batch size and data throughput. Lowering the resolution exacerbates information loss, leading to inferior results. To address this, Kong et al. [18] proposed Scalable Training from Range view (STR), a strategy that divides range images into multiple sub-images from different perspectives to reduce memory consumption. Although STR lessens memory usage, it results in a slight drop in segmentation accuracy and only minor improvement in inference speed compared to the baseline. Augmentation Data augmentation plays a crucial role in helping models learn more generalized representations, thereby enhancing scalability. For example, Mix3D [25] introduced an out-of-context mixing strategy by fusing two scenes. Similarly, MaskRange [13] proposed a weighted paste-drop augmentation to manually balance the class frequencies, while RangeFormer [18] employed four consecutive range-wise operations to provide richer semantic and structural cues in the scene. In this work, we introduce two novel augmentation steps, including point-wise and range-wise fusion, specifically designed for the range images utilized in our new schema. Post-Processing Addressing the prevalent ”many-to-one” problem in range-view representations often necessitates a post-processing step to upsample 2D predictions, a critical yet underexplored area in prior research. For example, RangeViT [1] introduced a trainable 3D refiner using KPConv [32]. However, while it directly optimizes 3D semantics, the performance improvement is limited, and the approach adds significant computational overhead. Some methods rely on conventional unsupervised techniques to infer semantics for 3D points; for instance, Milioto et al. [24] proposed a KNN-based voting approach, and Zhao et al. [48] introduced Nearest Label Assignment (NLA), which assigns labels based on the closest labeled point in 3D space. Nevertheless, these unsupervised techniques often struggle with accurately predicting boundaries and distant points, with performance further declining as range-image resolution decreases. To overcome these limitations, we design a new post-processing method that better interpolates the predictions of unprojected points using neighborhood information.
3 Proposed Approach
Our method introduces critical optimizations across three areas: a distinct training and inference scheme for low-resolution range images, advanced augmentation techniques, and an effective post-processing approach for accurately mapping 2D predictions into 3D space.
3.1 Pre-processing
Range-view Projection A LiDAR point cloud consists of points captured during a single revolution, denoted as $P=\{p_{1},...,p_{n}\}$, where each measurement represents a 4D point including the cartesian coordinates $p_{i}=\{x_{i},y_{i},z_{i}\}$ and intensity $t_{i}$. By pre-defining the 2D resolution, assuming $W$ and $H$ as the image width and height, we can project the point cloud into a range image, where each row $v$ and column $u$ correspond to the elevation and azimuth angles of LiDAR points. The mathematical expression of the spherical projection model is as following:
$\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}\frac{W}{2}-\frac{W}{2\pi}arctan(\frac{y}{x})\\ \frac{H}{\Theta_{max}-\Theta_{min}}*(\Theta_{max}-arcsin(\frac{z}{d}))\end{bmatrix}$ (1)
Angular values $\Theta_{max}$ and $\Theta_{min}$ define the upper and lower bound of the LiDAR’s vertical field of views and the depth value is calculated by $d=\sqrt{x^{2}+y^{2}+z^{2}}$. Note that $H$ is typically determined by the number of LiDAR sensor beams, while W can be assigned with random value based on the requirements. Similar to the prior studies [5, 48, 7], we adopt a five-channel input representation ($x,y,z,t,d$). Data Representation Increasing azimuth resolution can help preserve more details from point clouds, but this comes at the expense of reduced efficiency. Additionally, as illustrated in Fig. 2, elevation resolution plays an equally important role in maximizing projection rates. This led us to rethink if a large image width is necessary to mitigate information loss. Inspired by STR [18], we propose an alternative solution: downsampling the point cloud into multiple equal-interval sub-clouds and projecting them into range images with a lower azimuth resolution. During training, we randomly select one range image. For inference, we stack all images along the batch dimension and process them in a single forward pass. We depict the workflow of FLARES in Fig. 3. To balance 3D validity and 2D occupancy, assuming $N_{max}$ is the maximum number of partitioning groups for the specific resolution of range images, it is determined by the rule that the average 2D occupancy must not fall below the high-resolution range image in standard mode ($\frac{1}{N_{max}}\sum^{N_{max}}_{i}Occ_{i}\geq Occ_{high}$). This new design offers three key advantages: 1) Enhanced projection rate by increasing both image height and width; 2) Reduced memory consumption, enabling deployment on smaller GPUs; 3) Preservation of the full field of view, maintaining contextual integrity despite downsampling.
Figure 3: Visual Illustration of FLARES: The full LiDAR point cloud is equally divided and grouped, each projected into a lower-resolution range image. During training, one image and its 2D label are sampled for optimization. For testing, stacked sub-cloud projections are processed simultaneously, and the outputs are fused into 3D predictions using an unsupervised method. We provide more details in the supplementary material.
3.2 Data Augmentation
Previous studies [5, 1, 48] have primarily used geometric transformations such as random flipping, translation, and rotation for point cloud augmentation. To further enrich semantic contexts and optimize for the low-resolution range image used in FLARES, we introduce two additional augmentation steps: 1) Weighted Paste-Drop+ Class imbalance is a common issue in LiDAR semantic segmentation benchmarks [2, 3], where certain classes are heavily underrepresented. This imbalance is compounded by information loss during point cloud-to-range image conversion, especially when downscaling azimuth resolution. Building upon the Weighted Paste Drop (WPD) approach introduced in MaskRange [13], which pastes pixels from rare classes and drops pixels from abundant classes, we present an enhanced version, WPD+. Unlike the original method, which performs geometric data augmentation identically on both sampled and current frames in 3D space before projection, our approach applies WPD directly in 3D space, which avoids the repeating computation of geometric transformations, and samples multiple frames to improve class balancing. Additionally, we use a small set of synthetic dataset generated in the Carla Simulator [10] to further augment rare classes that correspond to small and dynamic objects in the scene. Despite possible domain gaps between datasets, it yields notable accuracy improvements from our experimental results. 2) Multi-Cloud Fusion Given the inherent sparsity of LiDAR point clouds, range images often contain a considerable number of empty grids. Prior works have mitigated the issue by interpolating missing pixels from nearest neighbors [44] or by supplementing occupied pixels from other frames [18]. Coming down to the point cloud splitting in FLARES mode, the 2D occupancy in the single range image is decreased and it prompts the necessity for some solution. In our empirical study in Fig. 5(d), we found that it yields sub-optimal results when training directly on sub-cloud. To address the issue, we propose Multi-Cloud Fusion (MCF), a strategy to increase the 2D occupancy of sub-clouds alongside the setup of FLARES. Assuming the point cloud can be divided into a maximum of $N_{max}$ groups, we randomly pick a group number $N$in $\{1,2,...,N_{max}\}$ and split the full point cloud accordingly. After projecting them into $N$ range images, we randomly select one $R\in\mathbb{R}^{H\times W\times 5}$ as the training input. To enhance occupancy, all empty pixels in $R$ are filled using occupied pixels from remaining $N-1$ range images. This method maximizes the 2D occupancy in the range image of a sub-cloud while maintaining the structural consistency of the scene.
For further technical details of how input data is curated and augmented, please refer to the supplementary material.
3.3 Post-Processing
After the augmented image being processed, 2D predictions from the network must be reprojected into 3D space using some post-processing technique. To align with the new inference framework with stacked predictions, we first propose an extension of the standard KNN method [24], termed  KNN Ensembling. In the post-processing phase, all sub-clouds are iteratively processed with KNN and votes are ensembled for every point from the full cloud to obtain final predictions. However, we found that the extension still faces the challenge in appropriately weighting contributions of nearest neighbors in 3D coordinates. In addition, the inference time is accumulated due to iterative process. To address the limitation, we propose a novel algorithm called Nearest Neighbors Range Interpolation (NNRI). Its pseudo-code is detailed in Algo. 1.
After applying softmax to the network output, we begin by kernelizing 2D predictions and range images using a pre-defined kernel size ($3\times 3$ in our experiments). Next, we assign each point’s nearest neighbors in 2D space with corresponding 2D coordinates and stack them along the sub-cloud dimension. The relative depth between each point and its neighbors is computed by taking the absolute difference in depth values. To extract valid data for interpolation, a threshold is needed to filter out distant neighbors. According to the prior knowledge [16, 20], using a constant threshold is sub-optimal due to differing point densities in LiDAR data: closer points are more likely to be affected by outliers due to high density, while farther points struggle to find valid neighbors due to sparsity. To fit this underlying geometry, the range value of each point is employed to determine its cut-off value. By normalizing the range using pre-computed mean and standard deviation, the cut-off value is derived from an exponential function, which approximates the relationship between point-sensor distance and density [21]. This approach simplifies computation by avoiding the costly nearest neighbor search in 3D space to calculate exact density values and adaptively assigns a threshold to each point. Once valid nearest neighbors are identified, they are normalized within the range of $[0,1]$ to compute interpolation weights. Finally, softmax scores of all 3D points are interpolated by the weighted sum of their nearest neighbors. NNRI is designed to effectively mitigate the ”many-to-one” issue inherently in range-view methods by leveraging distance-wise local neighborhood information in both 2D and 3D.
Define : $N=N_{max}$ sub-clouds.
The annotation contains $C$ classes
Input : Range images $R_{ranges}$ with size $N\times H\times W$,
Softmax scores $I_{scores}$ with size $N\times C\times H\times W$,
Arrays $R_{all}(p)$ with range values for all points,
Image coordinates $(u_{all},v_{all})$ for all points,
Kernel size $k$,
Padding $pad$,
Cut-off factor $\alpha$,
Mean of all range values $r_{mean}$,
Standard Deviation of all range values $r_{std}$
Output : Array $Labels$ with predicted labels for all points.
1:Unfold scores and ranges with $k\times k$ kernel:
2:    $S_{s}(n,h,w,k)\leftarrow\texttt{unfold}(I_{scores},k,pad)$
3:    $S_{r}(n,h,w,k)\leftarrow\texttt{unfold}(R_{ranges},k,pad)$
4:Extract nearest-neighbors for each point $p$:
5:    $N_{s}(n,p,k)\leftarrow S_{s}(n,h,w,k)[...,u_{all},v_{all}]$
6:    $N_{r}(n,p,k)\leftarrow S_{r}(n,h,w,k)[...,u_{all},v_{all}]$
7:Compute relative depths:
8:    $N_{rel}(n,p,k)\leftarrow||(N_{r}(n,p,k)-R_{all}(p)||$
9:Compute the cut-off value for each point $p$:
10:    $D(p)=\texttt{exp}(\frac{R(p)-r_{mean}}{r_{std}})*\alpha$
11:Filter the valid neighbors and compute weights:
12: $N_{valid}(n,p,k)\leftarrow\texttt{clamp}(N_{rel}(n,p,k),\texttt{max}=D(p)$)
13:    $W(n,p,k)=1-\texttt{Normalize}(N_{valid}(n,p,k))$
14:Weighted Sum for 3D Projection:
15:    $Scores(p)=\sum_{i}^{k^{2}\times n}W(n,p,k)*N_{s}(n,p,k)$
16:    $Labels=\texttt{argmax}_{c\in C}(Scores(p))$
17:Return $Labels$
Algorithm 1 Nearest Neighbors Range Interpolation
3.4 Network Selection
In order to pursue enhancement in the segmentation accuracy while possibly maintaining the high efficiency of range-view-based approaches, we revisited prior works and selected three light-weight CNN-based networks (FIDNet [48], SalsaNext [7] and CENet [5]) for integration. To further test the effectiveness across different network architectures, we additionally deploy RangeViT [1], a network composed of a series of Vision Transformer blocks [9], in the experimental phase. Original RangeViT uses a trainable KPConv-based 3D projector to get the point-wise predictions. We replace it with our post-processing component to achieve the full integration of our framework and train the model from scratch.
4 Experimental Analysis
4.1 Settings
Datasets We conduct experiments on two public LiDAR semantic segmentation datasets. SemanticKITTI [2] dataset [2] consists of 22 sequences captured with a 64-beam LiDAR sensor, encompassing 19 semantic classes. The dataset is split as follows: sequences 00 to 10 (excluding 08) are used for training, sequence 08 is reserved for validation, and sequences 11 to 21 are designated for testing. nuScenes dataset [3] comprises 1,000 driving scenes recorded in Boston and Singapore using a 32-beam LiDAR sensor, leading to a relatively sparse point cloud. After merging similar and infrequent classes, the dataset includes 16 distinct semantic classes. Implementation Details Prior works experimented mostly with the resolution of $64\times 2048$ for SemanticKITTI [5, 24, 7], and $32\times 960$ [18] or $32\times 2048$ [1] for nuScenes. In contrast, we reduce the azimuth resolution while increasing the projection rate in FLARES mode: resolutions of $64\times 512$ for SemanticKITTI and $32\times 480$ for nuScenes are fixed for the input and the full point cloud is split into up to 3 and 2 ($N_{max}$) sub-clouds during training and inference. We directly use custom configurations of prior works [48, 5, 7, 1] to train the networks in FLARES mode. For training the selected models (excluding RangeViT) on the nuScenes dataset, we standardize the hyperparameter set since no default configurations are provided. Specifically, we use the AdamW optimizer [22] along with a OneCycle scheduler [29], setting the maximum learning rate to $1e^{-3}$ and training for 150 epochs. All models are trained on four NVIDIA GeForce GTX 1080Ti in distributed mode. Evaluation Metrics Following prior works, we assess the performance using Intersection-over-Union (IoU) $\text{IoU}_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ and Accuracy (Acc) $\text{Acc}_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}}$ for each class $i$, and calculate the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) across all classes. $TP_{i}$, $FP_{i}$, and $FN_{i}$ represent the true positives, false positives, and false negatives for class $i$, respectively.
4.2 Comparative Study
Groundtruth Baseline FLARES
SalsaNext [7]
Refer to caption Refer to caption Refer to caption
FIDNet [48]
Refer to caption Refer to caption Refer to caption
CENet [5]
Refer to caption Refer to caption Refer to caption
RangeViT [1]
Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative results on SemanticKITTI[2] Points in red and gray represent incorrect and correct predictions, respectively. More examples are provided in the supplementary material.
We compare FLARES with baseline models across two datasets. As shown in Tab. 1, all four networks see significant improvements: SalsaNext has mIoU gains of 5.3% on SemanticKITTI and 2.3% on nuScenes, FIDNet improves by 7.9% and 3.9%, and CENet by 3.3% and 3.1%. RangeViT, as a ViT-based network, also exhibits huge enhancement in performance, confirming FLARES’s generalizability across different architectures. This improvement is especially prominent for smaller, dynamic, and under-represented classes such as truck, motorcycle, bicycle, pedestrian and bicyclist. In Fig. 4, we present a visual comparison. Notably, with the support of FLARES, the network demonstrates improved accuracy in segmenting foreground objects. An exception arises with the motorcyclist class in SemanticKITTI, where IoU scores decrease compared to the baseline. Diving into the problem, this can be traced back to the extremely low occurrence of annotations for that class in the dataset. In standard training on low-resolution range images, this class already suffers from poor representation. In FLARES mode, the occurrence is further reduced by splitting of the point cloud. This accumulation of downsampling prevents the network from optimizing on that rare class effectively and lead to inferior performance. In contrast, the improvement on nuScenes is more consistent as class frequencies are better balanced. We regard this as a corner case when testing on an class-imbalanced dataset. As a future work to resolve the issue, we aim to explore 3D reconstruction techniques to generate real-world-like pseudo LiDAR point clouds for augmentation [4, 23].
In Tab. 2, we further compare FLARES-boosted networks with other state-of-the-arts from various modalities. Given that methods we select preserve relatively fewer number of parameters, FLARES helps to improve the segmentation accuracy, being comparable to other point- or voxel-based approaches that deploy much larger and deeper networks. In addition, they outperform others significantly in latency, achieving superior trade-offs in accuracy and efficiency.
 SemanticKITTI test set
Method
mIoU
car
bicy
moto
truc
o.veh
ped
b.list
m.list
road
park
walk
o.gro
build
fenc
veg
trun
terr
pole
sign
SalsaNext [7] $59.5$ $91.9$ $48.3$ $38.6$ $38.9$ $31.9$ $60.2$ $59.0$ $\mathbf{19.4}$ $91.7$ $63.7$ $75.8$ $\mathbf{29.1}$ $90.2$ $64.2$ $81.8$ $63.6$ $\mathbf{66.5}$ $54.3$ $62.1$
[…]
FIDNet [48] $51.3$ $90.4$ $28.6$ $30.9$ $34.3$ $27.0$ $43.9$ $48.9$ $16.8$ $90.1$ $58.7$ $71.4$ $19.9$ $84.2$ $51.2$ $78.2$ $51.9$ $64.5$ $32.7$ $50.3$
FIDNet $59.5$ $93.9$ $54.7$ $48.9$ $27.6$ $23.9$ $62.3$ $59.8$ $\mathbf{23.7}$ $90.6$ $59.1$ $75.8$ $26.7$ $88.9$ $60.5$ $84.5$ $64.4$ $69.0$ $53.3$ $62.8$
[…]
⋄†CENet [5] $60.7$ $92.1$ $45.4$ $42.9$ $43.9$ $46.8$ $56.4$ $63.8$ $29.7$ $91.3$ $66.0$ $75.3$ $31.1$ $88.9$ $60.4$ $81.9$ $60.5$ $67.6$ $49.5$ $59.1$
CENet $64.7$ $91.9$ $58.6$ $50.3$ $40.6$ $42.3$ $68.9$ $65.9$ $\mathbf{43.5}$ $90.3$ $60.9$ $75.1$ $31.5$ $91.0$ $66.2$ $84.5$ $69.7$ $\mathbf{70.0}$ $61.5$ $67.6$
[…]
RangeViT [1] $64.0$ $95.4$ $55.8$ $43.5$ $29.8$ $42.1$ $63.9$ $58.2$ $\mathbf{38.1}$ $\mathbf{93.1}$ $70.2$ $\mathbf{80.0}$ $32.5$ $\mathbf{92.0}$ $\mathbf{69.0}$ $\mathbf{85.3}$ $\mathbf{70.6}$ $\mathbf{71.2}$ $\mathbf{60.8}$ $64.7$
[…]
 nuScenes val set
Method (year)
mIoU
barrier
bicy
bus
car
const
moto
ped
traffic.c
trailer
truck
driv
o.flat
side
terrain
manm
veg
SalsaNext $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $\mathbf{63.1}$ $61.3$ $76.5$ $96.0$ $\mathbf{70.8}$ $71.2$ $71.5$ $86.7$ $84.4$
[…]
FIDNet $72.7$ $73.0$ $36.0$ $87.8$ $86.0$ $45.6$ $74.1$ $73.9$ $62.5$ $67.1$ $77.7$ $94.3$ $69.8$ $72.2$ $72.1$ $86.1$ $84.5$
[…]
CENet $73.7$ $73.6$ $32.9$ $92.7$ $87.1$ $\mathbf{53.5}$ $76.1$ $69.0$ $58.7$ $66.8$ $81.6$ $95.6$ $71.1$ $73.7$ $73.2$ $87.5$ $85.7$
[…]
RangeViT $75.2$ $75.5$ $\mathbf{40.7}$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $\mathbf{66.3}$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $\mathbf{89.9}$ $\mathbf{87.2}$
[…]
Table 1: Comparisons of state-of-the-art LiDAR semantic segmentation methods on the test set of SemanticKITTI [2] and val set of nuScenes [3] in standard and  FLARES mode. IoU scores are reported in percentages (%). For each method block, bold and underline indicate the best and second best result in the column. Baseline results trained on low-resolution ($64\times 512$) range images. Models inferred with test-time augmentation [18]. Note that we did not use model ensembling to further boost the model performance.
[…]
PolarNet [46] [’20] 13.6M 71 Polar 71.0 54.9 57.5
SPVNAS [31] [’20] 12.5M 259 Voxel - 64.7 66.4
RandLA-Net [42] [’20] 1.2M 55 Point - - 53.9
Tornado-Net [12] [’20] - - Multiple - 64.5 63.1
[…]
Cylinder3D [49] [’21] 56.3M 170 Voxel 76.1 67.8 65.9
RPVnet [43] [’21] 24.8M 168 Multiple 77.6 68.2 70.3
FPS-Net [40] [’21] 55.7M 48 Range - 54.9 57.1
Lite-HDSeg [28] [’21] - 50 Range - 64.4 63.8
[…]
Meta-RSeg [34] [’22] 6.8M 46 Range - 60.3 61.0
PVKD [15] [’22] 14.1M 76 Voxel 76.0 66.4 71.2
PTv2 [38] [’22] 12.8M 213 Point 80.2 70.3 72.6
2DPASS [45] [’22] 26.5M 119 Multiple 79.4 69.3 72.2
GFNet [27] [’22] - 100 Multiple 76.8 63.2 65.4
[…]
4.3 Ablation Study
To perform the ablation study, we test with CENet [5] on val set of SemanticKITTI [2]. Training Schema As shown in Tab.3, we conduct a comprehensive evaluation to examine the trade-off between accuracy and efficiency under different training configurations. In the standard mode, a high-resolution range image is used as the input to reproduce baseline results. While training in the STR paradigm[18], there is a slight performance drop, the memory consumption is significantly reduced by partitioning the full point cloud through azimuth-wise grouping. The limited increase in latency is due to the fact that STR has to additionally unite the batched prediction along azimuth resolution and the post-processing is still performed on the high-resolution image. In contrast, FLARES achieves remarkable improvements, showing increases of 2.7% in mIoU, 2.1% in mAcc, and 45% acceleration in inference. Similar to STR, FLARES reduces memory consumption compared to the standard mode, further optimizing the efficiency of range-view semantic segmentation.
Method Input resolution mIoU mAcc Lat.
Standard $1\times 64\times 2048$ 64.8 77.2 44 ms
STR [18] $1(5)\times 64\times 480$ 64.3 76.8 41 ms
FLARES $1(3)\times 64\times 512$ 67.5 79.3 24 ms
Table 3: Performance with varying training schemes is evaluated on val set of SemanticKITTI [2]. Input resolution is formatted in $B\times H\times W$ (batch size, image width and height). For STR [18], we use the configuration yielding the best performance on CENet [5]: a full resolution of $64\times 1920$ split into four $64\times 480$ sub-images. For the fair comparison, all models are integrated with proposed components and trained from scratch.
Data Augmentation As presented in Fig. 5(a), various data augmentation methods can enhance the segmentation performance by large margins and our WPD+ performs the best among them, which increases baseline mIoU by 5.7% and mAcc by 2.9%. Mix3D [25] increases the contextual information per frame by fusing one scene to another, however, the model can still suffer from class imbalance. RangeAug [18] consists of 4 different range-wise operations to enrich the semantic and structural cues. From their experiments, it demonstrated that this augmentation technique is especially effective for attention-based networks [33, 9], which possess significantly higher model capacities and are more reliant on data diversity for optimal performance. Similar to their experimental results, applying this method to lightweight CNN-based networks has shown limited success in improvement.
(a) WPD+
(b) #Frames sampled in WPD+
(c) Inclusion of synthetic dataset
(d) MCF
Figure 5: a) Initialization of training with standard geometric data augmentation (GDA) and benchmark several state-of-the-art 3D augmentation techniques, including Mix3D [25], RangeAug [18], and original WPD [13]. b) Different number of sampled frames for best performance c) A comparative plot showing the IoU scores of top-rare classes in scenarios both with and without the inclusion of the synthetic dataset. As reference, the class frequencies in the val set are provided. d) The models are trained using two different input configurations: either a single range image derived from the full cloud (FC) or a range image generated from a sub-cloud (SC), and inferred in FLARES mode. Note that for all trained models in a) - d), we leverage KNN Ensembling during post-processing phase.
WPD+ includes two tunable parameters: the number of sampled frames from the original dataset and the use of the synthetic dataset. To find out the best parameter set, we conduct two additional experiments. Fig. 5(b) shows that sampling 6 frames results in the optimal performance, while increasing the number of frames beyond this point leads to performance degradation. This is likely due to the model’s limited scalability, similar to the effects observed in case of RangeAug[18], in addition, pasting too many pixels belonging to specific classes will again corrupt the semantic balance. Furthermore, Fig. 5(c) illustrates that the synthetic dataset plays a key-role in refining semantic prediction of top-rare classes. Noteworthily, using the synthetic dataset is efficient and practical because this allows us to customize sensor configurations to align with the target dataset and to define specific objects within the scene for downstream applications without any labor cost.
In the next stage, we study how unoccupied pixels in the range image can affect the performance. Firstly, we use the full point cloud for training and sub-clouds for inference, as the decreasing azimuth resolution can already resolve the problem of low 2D occupancy (the first column in Fig. 5(d)). However, empirical results reveal a drop in 2D accuracy when inferring from sub-clouds. This is possibly because of the domain shift caused by the occupancy difference in range images. In the next trial, we use the sub-cloud as training input instead, but this directly limits the performance due to the occupancy reduction. To tackle this compounded challenge brought by FLARES, we introduce Multi-Cloud Fusion, an additional data augmentation step which fuses multiple sub-clouds through occupancy padding during projection phase. As shown in Fig. 5(d), using MCF exhibits the highest performance in the series, achieving 1% increase in IoU over the model trained on full cloud. As an alternative, we apply RangeIP [44], an interpolation-based augmentation technique, to enhance 2D occupancy in the range image, but it results in slight worse accuracy compared to the baseline.
(a) Ablation on Post-Processing
(b) Inference Time [ms]
(c) Ablation on Input Resolution
Figure 6: a) - b): Various post-processing techniques are applied to the same trained model. KNN [24] and NLA [48] are used on the single range image generated from the full point cloud, whereas KNN Ensembling operates on multiple range images derived from sub-clouds. For NNRI, we assess it using both full cloud and sub-clouds. c): Evaluation on different input resolutions and corresponding number of sub-clouds ($N_{max}$).
Post-Processing As introduced in Sec. 3.3, a simple LiDAR model is built to approximate the distance-density function of 3D points and compute cut-off values for valid neighbors extraction. To verify the necessity of the step, some qualitative results are provided in Fig. 7. As can be seen, the adaptive cut-off values refine semantic predictions by better accommodating objects with varying density scales.
(a) Scene
(b) GT
(c) Constant
(d) Adaptive
Figure 7: Segmentation results with different cut-off values in NNRI: in the case of constant value (set at 1), overlapping points of Road are partially misclassified as Car in the top image. Similarly, in the bottom image, half of the points that belong to Building are incorrectly predicted as Vegetation.
Next, we explore the impact of various post-processing techniques on segmentation performance in Fig. 6(a) and 6(b). Regarding conventional KNN [24] as the baseline, NLA [48] demonstrates similar performance in both accuracy and latency. In contrast, we deploy our approach (NNRI) in the standard mode as well and observe a significant improvement: inference time is cut nearly 16% compared to KNN, while mAcc and mIoU increase by 2.8% and 2%, respectively. Unlike KNN, NNRI avoids the computational cost of Gaussian kernel calculations for distance weighting and directly performs nearest neighbor searches on the range image instead of in 3D space, further reducing computational overhead. NNRI interpolates class-wise scores based on relative depths rather than directly voting on hard labels, relying more on weighted information from nearest neighbors, which is the major reason why it outperform other post-processing approaches.
Switching to FLARES mode, we first extend the standard KNN approach to KNN Ensembling, which iteratively gathers votes from all points in each sub-cloud and aggregates them for the final prediction. While this extension improves the accuracy, it comes at approximately doubled latency cost. Conversely, when NNRI is adapted, it consistently provides notable improvements in both efficacy and efficiency. As a reference, we included evaluation scores on 2D predictions (-- dashed lines in Fig. 6(a)), showing that FLARES with NNRI significantly narrows the accuracy gap between 2D and 3D predictions. This suggests that our approach effectively mitigates the ”many-to-one” problem, offering substantial gains in segmentation performance. Input Resolution To further explore the optimal results in FLARES mode, we test various input resolutions, as shown in Fig. 6(c). From the experimental results, we found that resolutions of $512$ and $640$ deliver the best mIoU scores. Increasing the azimuth resolution beyond this point causes a slight performance drop. Nevertheless, all tested configurations outperform the baseline (the first row in Tab.3), demonstrating the superb effectiveness of our approach on range-view LiDAR semantic segmentation.
5 Conclusion
In this work, we introduced FLARES, an optimized training and inference schema designed for seamless integration into any range-view-based network. To enhance the scalability, we further developed two data augmentation techniques for class rebalance and occupancy increase. Additionally, we proposed a novel unsupervised post-processing method to effectively address the ”many-to-one” issue, minimizing the 2D-to-3D performance gap. Our approach has substantiated significant improvements in accuracy and efficiency over baselines across various network architectures on two widely used LiDAR benchmarks. Despite the limitation in some corner case of class imbalance in the dataset, our approach shows its overall versatility and effectiveness in advancing LiDAR semantic segmentation.

### v2

@@
-Bin Yang1,2, Alexandru Paul Condurache1,2
+ΥFLARESΥ: Fast and Accurate LiDAR Multi-Range Semantic Segmentation
 Abstract
-3D scene understanding is a critical yet challenging task in autonomous driving, primarily due to the irregularity and sparsity of LiDAR data, as well as the computational demands of processing large-scale point clouds.
-Recent methods leverage the range-view representation to improve processing efficiency.
-To mitigate the performance drop caused by information loss inherent to the ”many-to-one” problem, where multiple nearby 3D points are mapped to the same 2D grids and only the closest is retained, prior works tend to choose a higher azimuth resolution for range-view projection.
-However, this can bring the drawback of reducing the proportion of pixels that carry information and heavier computation within the network.
-We argue that it is not the optimal solution and show that, in contrast, decreasing the resolution is more advantageous in both efficiency and accuracy.
-In this work, we present a comprehensive re-design of the workflow for range-view-based LiDAR semantic segmentation.
-Our approach addresses data representation, augmentation, and post-processing methods for improvements.
-Through extensive experiments on two public datasets, we demonstrate that our pipeline significantly enhances the performance of various network architectures over their baselines, paving the way for more effective LiDAR-based perception in autonomous systems.
-The code will be released based on the acceptance.
+3D scene understanding is a critical yet challenging task in autonomous driving due to the irregularity and sparsity of LiDAR data, as well as the computational demands of processing large-scale point clouds.
+Recent methods leverage range-view representations to enhance efficiency, but they often adopt higher azimuth resolutions to mitigate information loss during spherical projection, where only the closest point is retained for each 2D grid.
+However, processing wide panoramic range-view images remains inefficient and may introduce additional distortions.
+Our empirical analysis shows that training with multiple range images, obtained from splitting the full point cloud, improves both segmentation accuracy and computational efficiency.
+However, this approach also poses new challenges of exacerbated class imbalance and increase in projection artifacts.
+To address these, we introduce FLARES, a novel training paradigm that incorporates two tailored data augmentation techniques and a specialized post-processing method designed for multi-range settings.
+Extensive experiments demonstrate that FLARES is highly generalizable across different architectures, yielding 2.1%–7.9% mIoU improvements on SemanticKITTI and 1.8%–3.9% mIoU on nuScenes, while delivering over 40% speed-up in inference.
+[^1] Project page: https://binyang97.github.io/FLARES
 1 Introduction
@@
 In general, those methods can be categorized based on LiDAR data representation into point-based [64, 38, 53], voxel-based [66, 20] and projection-based methods [63, 34, 7, 65].
-Both point- and voxel-based approaches typically require substantial computational resources due to the need to process data through networks with numerous 3D convolutional layers, intensive feature pre-processing, and deep architectures involving multiple downsampling and upsampling operations.
-These requirements can result in slow inference speeds, limiting their suitability for real-time applications.
+Both point and voxel-based approaches typically require substantial computational resources due to the need to process data through networks with numerous 3D convolutional layers, intensive feature pre-processing, and deep architectures involving multiple downsampling and upsampling operations.
+These requirements can result in slow inference speeds, limiting their suitability for real-time applications [31].
 In contrast, rasterizing point cloud into range-view images [10] is more advantageous in fast and scalable LiDAR perception, because it allows the use of 2D operators for efficient computation and facilitates the transfer of knowledge from camera images [2, 25].
+(a) Standard
+(b) STR [25]
+(c) FLARES
 […]
-Nevertheless, learning from range-view representations can suffer from the performance drop caused by the ”many-to-one” conflict of adjacent points.
-To resolve the problem, most approaches maximize the azimuth resolution to make the range image more informative [24, 7, 5, 48, 13, 18].
-We argue that this is suboptimal for the following reasons: 1) High overhead: increasing resolution significantly raises computational demand, limiting real-time performance.
-2) Inefficient computation: the sparse nature of LiDAR data leads to many empty grid cells.
-Those unoccupied pixels introduce noise in the data and consume resources needlessly.
-3) Limited informative gains: due to distortions caused by ego-motion and sensor-internal errors, points mapped to the same azimuthal location do not perfectly align with the laser beams.
-This misalignment implies that prioritizing both azimuth and elevation resolutions, rather than increasing image width alone, would yield a higher projection rate, as seen in our analysis in Fig.
-2.
+Nevertheless, learning from range-view representations can suffer from a ”many-to-one” conflict among adjacent points.
+This issue causes an irreversible loss of spatial information and leads to inferior performance compared to other methods that utilize 3D LiDAR representations.
+Although increasing the azimuth resolution of range images can alleviate this loss, it forces the network to process wide panoramic images (Fig.
+1(a)), which significantly increases computational overhead [34, 10, 7, 65, 19, 25].
+Building on this standard setting, recent work introduces a scalable training framework with range-view images (STR) to improve efficiency [25].
+As shown in Fig.
+1(b), by splitting the point clouds into multiple distinct views, STR processes all crops in a batch during inference.
+The resulting predictions are then merged to reconstruct the full range image, which is subsequently post-processed to generate 3D coordinates.
+However, the splitting strategy of STR is heuristic and compromises scene completeness.
+Empirical results in [25] reveal a slight performance drop within some networks.
+Furthermore, the gain in inference speed is limited when maintaining the high resolution of range images during the projection and post-processing stages.
+To solve these challenges, we propose FLARES, a novel training paradigm for range-view LiDAR semantic segmentation.
+Its procedure is visually illustrated in Fig.
+1(c).
+In contrast to cartesian grid splits of STR [25], FLARES divides point clouds along the LiDAR’s spherical coordinate and then projects each split into lower-resolution range images.
+This design preserves local geometric coherence and scene integrity during partitioning, which enables more effective semantic context sharing across generated range images compared to STR.
+Furthermore, our method exclusively processes low-resolution images, achieving greater efficiency than both STR and standard settings.
 Figure 2: Statistics on SemanticKITTI [3]: 3D validity (proportion of projected points) with different azimuth (W) and elevation (H) resolutions.
 Comparable increases are observable when doubling azimuth and elevation resolution ($\Delta V_{azi},\Delta V_{ele}$).
-Building on the aforementioned insights, we introduce FLARES, a novel training schema for range-view semantic segmentation of LiDAR point clouds.
-Rather than focusing on designing a new network architecture, FLARES targets to address the common limitations outlined above, hence, our proposal is generalizable to any range-view-based approaches.
-In essence, FLARES divides the point cloud into multiple sub-clouds, with only one projected onto a low-resolution image during training.
-During inference, all sub-clouds are projected and stacked along the batch dimension for processing.
-This approach increases the projection rate in both azimuth and elevation, thereby enriching informativeness of the range-view representation at lower cost.
-Downscaling resolution, however, can exacerbate class imbalance in the dataset [7], potentially leading to overfitting during training.
-Another problem is decreasing 2D occupancy due to splitting the point cloud.
-To tackle these two incidental issues, we extend the pipeline with two additional data augmentation steps.
-Furthermore, we explore improvements in post-processing methods, a topic that has received minimal attention in previous works.
-In summary, our contributions are as follows: 1) We introduce FLARES, a newly designed training schema for faster and more accurate LiDAR semantic segmentation.
-2) We integrate two data augmentation techniques tailored to the new schema to enhance the network performance.
-3) We propose a novel interpolation-based post-processing approach to resolve the ”many-to-one” problem more effectively.
-4) We generalize this schema and all components across various range-view semantic segmentation networks, demonstrating superior performance over baselines on two different benchmarks.
+Moreover, multi low-resolution range-view images created by FLARES are generally a more informative and powerful representation compared to the high-resolution projection.
+Due to ego-motion and sensor calibration error, points mapped to identical azimuthal positions often misalign with actual laser beam orientations [61].
+This inherent mismatch demonstrates that jointly prioritizing azimuth and elevation resolutions, rather than solely expanding image width, achieves higher projection fidelity.
+Statistically, as quantified in our occupancy measurement (Fig.
+2), doubling the elevation resolution achieves approximately 80% of the projection gain obtained by doubling the azimuth resolution in SemanticKITTI [3] dataset.
+This property indicates that our new representation can restore more 3D points, thereby alleviating the ”many-to-one” conflict among adjacent points more effectively.
+(a) Before splitting
+(b) After splitting
+Figure 3: An example from SemanticKITTI [3] dataset is visualized in top-down view and range-view.
+Three crops are magnified to specify three limitations brought by the splitting of point clouds.
+1) Exacerbated class imbalance: objects with small sizes, likewise low occurrences in annotations, tend to fade away in the range image after downsampling (e.g.
+pole in the left crop).
+2) Intensified noise: reduction in points results in increasing clutters (middle crop) that may disrupt the training stability.
+3) Deteriorated distortion: decrease in point density can introduce more projection artifacts, thus corrupting the sharpness of local geometry (e.g.
+blurry boundaries of car in the right crop).
+While FLARES offers significant advantages, it introduces three key challenges, as visualized in Fig.
+3.
+First, point cloud splitting exacerbates class imbalance issues [6], where classes representing small objects with inherently sparse distributions become further undersampled.
+Second, the splitting operation increases point cloud sparsity, causing substantial reductions in occupancy of range images.
+These unoccupied pixels act as noise that destabilize training convergence [25].
+Finally, the resulting projection artifacts degrade image sharpness, particularly at partition boundaries where geometric discontinuities emerge.
+To tackle these incidental issues, we extend the pipeline with two data augmentation steps.
+Furthermore, we investigate in exploring a more effective post-processing technique tailored to the multi-range settings to further enhance the advantages of our method.
+In summary, our contributions are:
+  1. 1.
+    FLARES, a scalable training paradigm that enhances LiDAR semantic segmentation with improved speed, accuracy, and architectural agnosticism.
+  2. 2.
+    Two specialized data augmentation techniques addressing exacerbated class imbalance and noise amplification within FLARES.
+  3. 3.
+    A novel interpolation-based post-processing method that refines range-view predictions using precise nearest-neighbor information.
+  4. 4.
+    Comprehensive experiments across four network architectures and two benchmarks demonstrating consistent superiority over baselines.
 2 Related Works
-Point- and Voxel-based methods Some recent works [26, 47, 38] use raw point cloud data as direct network input, eliminating the need for post-processing after prediction.
+Point and Voxel-based methods Some recent works [38, 64, 52] use raw point cloud data as direct network input, eliminating the need for post-processing after prediction.
 However, these methods often face high computational complexity and memory usage.
@@
 However, the quadratic computational complexity of self-attention mechanisms in ViTs introduces challenges in achieving an optimal balance between efficiency and accuracy.
-Training schema Highlighting inefficiencies stemming from the use of high-resolution range images, some methods have adopted compact networks, significantly reducing network capacity [48, 7, 5].
-Unfortunately, high-resolution range images still demand substantial memory, which restricts scalability in terms of batch size and data throughput.
-Lowering the resolution exacerbates information loss, leading to inferior results.
-To address this, Kong et al.
-[18] proposed Scalable Training from Range view (STR), a strategy that divides range images into multiple sub-images from different perspectives to reduce memory consumption.
-Although STR lessens memory usage, it results in a slight drop in segmentation accuracy and only minor improvement in inference speed compared to the baseline.
-Augmentation Data augmentation plays a crucial role in helping models learn more generalized representations, thereby enhancing scalability.
-For example, Mix3D [25] introduced an out-of-context mixing strategy by fusing two scenes.
-Similarly, MaskRange [13] proposed a weighted paste-drop augmentation to manually balance the class frequencies, while RangeFormer [18] employed four consecutive range-wise operations to provide richer semantic and structural cues in the scene.
-In this work, we introduce two novel augmentation steps, including point-wise and range-wise fusion, specifically designed for the range images utilized in our new schema.
-Post-Processing Addressing the prevalent ”many-to-one” problem in range-view representations often necessitates a post-processing step to upsample 2D predictions, a critical yet underexplored area in prior research.
-For example, RangeViT [1] introduced a trainable 3D refiner using KPConv [32].
+Training Paradigm While existing methods address inefficiencies in high-resolution range images by employing compact networks to reduce model capacity [65, 10, 7], these approaches still require substantial memory for processing high-resolution inputs, limiting scalability in batch size and data throughput.
+Reducing resolution mitigates memory demands but worsens spatial information loss, degrading segmentation accuracy.
+To resolve this trade-off, Kong et al.
+[25] introduced Scalable Training from Range-view (STR), which splits range images into sub-images from multiple perspectives to lower memory consumption.
+However, STR’s use of partial scene views during training inherently limits segmentation accuracy.
+Furthermore, the method does not optimize for computational efficiency, leaving runtime improvements unaddressed.
+Post-Processing Addressing the prevalent ”many-to-one” problem in range-view representations often necessitates a post-processing step to upsample 2D predictions, a critical yet under-explored area in prior research.
+RangeViT [2] introduced a trainable 3D refiner using KPConv [46].
 However, while it directly optimizes 3D semantics, the performance improvement is limited, and the approach adds significant computational overhead.
-Some methods rely on conventional unsupervised techniques to infer semantics for 3D points; for instance, Milioto et al.
+Some methods rely on conventional unsupervised techniques to infer semantics for 3D points.
+For instance, Milioto et al.
 [34] proposed a KNN-based voting approach, and Zhao et al.
@@
 Nevertheless, these unsupervised techniques often struggle with accurately predicting boundaries and distant points, with performance further declining as range-image resolution decreases.
-To overcome these limitations, we design a new post-processing method that better interpolates the predictions of unprojected points using neighborhood information.
-3 Proposed Approach
-Our method introduces critical optimizations across three areas: a distinct training and inference scheme for low-resolution range images, advanced augmentation techniques, and an effective post-processing approach for accurately mapping 2D predictions into 3D space.
-3.1 Pre-processing
-Range-view Projection A LiDAR point cloud consists of points captured during a single revolution, denoted as $P=\{p_{1},...,p_{n}\}$, where each measurement represents a 4D point including the cartesian coordinates $p_{i}=\{x_{i},y_{i},z_{i}\}$ and intensity $t_{i}$.
-By pre-defining the 2D resolution, assuming $W$ and $H$ as the image width and height, we can project the point cloud into a range image, where each row $v$ and column $u$ correspond to the elevation and azimuth angles of LiDAR points.
-The mathematical expression of the spherical projection model is as following:
-$\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}\frac{W}{2}-\frac{W}{2\pi}arctan(\frac{y}{x})\\ \frac{H}{\Theta_{max}-\Theta_{min}}*(\Theta_{max}-arcsin(\frac{z}{d}))\end{bmatrix}$ (1)
+To overcome these limitations, we design a new post-processing method that adaptively weights neighboring contributions for interpolating predictions in 3D space.
+3 Methodology
+3.1 Proposed Framework
+A LiDAR point cloud consists of points captured during a single revolution, denoted as $P=\{p_{1},...,p_{n}\}$, where each measurement represents a 4D point including the cartesian coordinates $p_{i}=\{x_{i},y_{i},z_{i}\}$ and intensity $t_{i}$.
+In semantic segmentation task, each point has an additional feature of class label, denoted as $c_{i}$, for training.
+Our splitting configuration operates as follows: given a predefined partition count $N$, we distribute points into sub-clouds by along the spherical coordinate of the scan pattern.
+Specifically, a sub-cloud is derived as $P_{i}=\{p_{j}|j\mod N=i-1\}$ for $i\in\{1,2,..N\}$.
+This modulo-based assignment ensures equal partitioning while preserving spatial coherence within each sub-cloud.
+For the spherical projection, we specify the function $f$ as following:
+$f(p)=\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}\frac{W}{2}-\frac{W}{2\pi}arctan(\frac{y}{x})\\ \frac{H}{\Theta_{max}-\Theta_{min}}*(\Theta_{max}-arcsin(\frac{z}{d}))\end{bmatrix}$ (1)
+, where $W$ and $H$ as the image width and height, while $v$ and $u$ correspond to the elevation and azimuth angles of LiDAR points.
 Angular values $\Theta_{max}$ and $\Theta_{min}$ define the upper and lower bound of the LiDAR’s vertical field of views and the depth value is calculated by $d=\sqrt{x^{2}+y^{2}+z^{2}}$.
 Note that $H$ is typically determined by the number of LiDAR sensor beams, while $W$ can be assigned with random value based on the requirements.
-Similar to the prior studies [5, 48, 7], we adopt a five-channel input representation ($x,y,z,t,d$).
-Data Representation Increasing azimuth resolution can help preserve more details from point clouds, but this comes at the expense of reduced efficiency.
-Additionally, as illustrated in Fig.
-2, elevation resolution plays an equally important role in maximizing projection rates.
-This led us to rethink if a large image width is necessary to mitigate information loss.
-Inspired by STR [18], we propose an alternative solution: downsampling the point cloud into multiple equal-interval sub-clouds and projecting them into range images with a lower azimuth resolution.
-During training, we randomly select one range image.
-For inference, we stack all images along the batch dimension and process them in a single forward pass.
-We depict the workflow of FLARES in Fig.
-3.
-To balance 3D validity and 2D occupancy, assuming $N_{max}$ is the maximum number of partitioning groups for the specific resolution of range images, it is determined by the rule that the average 2D occupancy must not fall below the high-resolution range image in standard mode ($\frac{1}{N_{max}}\sum^{N_{max}}_{i}Occ_{i}\geq Occ_{high}$).
-This new design offers three key advantages: 1) Enhanced projection rate by increasing both image height and width; 2) Reduced memory consumption, enabling deployment on smaller GPUs; 3) Preservation of the full field of view, maintaining contextual integrity despite downsampling.
-Figure 3: Visual Illustration of FLARES: The full LiDAR point cloud is equally divided and grouped, each projected into a lower-resolution range image.
-During training, one image and its 2D label are sampled for optimization.
-For testing, stacked sub-cloud projections are processed simultaneously, and the outputs are fused into 3D predictions using an unsupervised method.
-We provide more details in the supplementary material.
+Similar to prior studies [7, 65, 10], we adopt a five-channel input representation ($x,y,z,t,d$).
+Next, we project all sub-clouds into range images simultaneously $R_{i}=f(P_{i})$.
+During training, since our splitting strategy leverages shared semantic contexts across sub-clouds, we randomly select only one for network optimization.
+During inference, all images $R_{i}$ are stacked as a batch and processed by the network to extract the 2D predictions.
+Unlike STR [25], which wraps all images to the original resolution, we treat each sub-cloud as an independent complete scene and apply post-processing to low-resolution range images individually.
+Final 3D predictions are reconstructed by merging outputs from all sub-clouds.
+An discussed in Sec.
+1, FLARES offers three key advantages: 1) Enhanced restoration of original 3D information through multi-range projection, 2) Reduced memory consumption by decreasing the image resolution, enabling the scalable training on smaller GPUs, 3) Preservation of the full field of view, which maintains contextual integrity despite downsampling.
 3.2 Data Augmentation
-Previous studies [5, 1, 48] have primarily used geometric transformations such as random flipping, translation, and rotation for point cloud augmentation.
-To further enrich semantic contexts and optimize for the low-resolution range image used in FLARES, we introduce two additional augmentation steps: 1) Weighted Paste-Drop+ Class imbalance is a common issue in LiDAR semantic segmentation benchmarks [2, 3], where certain classes are heavily underrepresented.
-This imbalance is compounded by information loss during point cloud-to-range image conversion, especially when downscaling azimuth resolution.
-Building upon the Weighted Paste Drop (WPD) approach introduced in MaskRange [13], which pastes pixels from rare classes and drops pixels from abundant classes, we present an enhanced version, WPD+.
-Unlike the original method, which performs geometric data augmentation identically on both sampled and current frames in 3D space before projection, our approach applies WPD directly in 3D space, which avoids the repeating computation of geometric transformations, and samples multiple frames to improve class balancing.
-Additionally, we use a small set of synthetic dataset generated in the Carla Simulator [10] to further augment rare classes that correspond to small and dynamic objects in the scene.
-Despite possible domain gaps between datasets, it yields notable accuracy improvements from our experimental results.
-2) Multi-Cloud Fusion Given the inherent sparsity of LiDAR point clouds, range images often contain a considerable number of empty grids.
-Prior works have mitigated the issue by interpolating missing pixels from nearest neighbors [44] or by supplementing occupied pixels from other frames [18].
-Coming down to the point cloud splitting in FLARES mode, the 2D occupancy in the single range image is decreased and it prompts the necessity for some solution.
-In our empirical study in Fig.
-5(d), we found that it yields sub-optimal results when training directly on sub-cloud.
-To address the issue, we propose Multi-Cloud Fusion (MCF), a strategy to increase the 2D occupancy of sub-clouds alongside the setup of FLARES.
-Assuming the point cloud can be divided into a maximum of $N_{max}$ groups, we randomly pick a group number $N$in $\{1,2,...,N_{max}\}$ and split the full point cloud accordingly.
-After projecting them into $N$ range images, we randomly select one $R\in\mathbb{R}^{H\times W\times 5}$ as the training input.
-To enhance occupancy, all empty pixels in $R$ are filled using occupied pixels from remaining $N-1$ range images.
-This method maximizes the 2D occupancy in the range image of a sub-cloud while maintaining the structural consistency of the scene.
-For further technical details of how input data is curated and augmented, please refer to the supplementary material.
+In particular, the point cloud splitting strategy of FLARES implicitly introduces two new challenges: exacerbated class imbalance and amplified projection noise.
+To address class imbalance, prior works [19, 15, 57] have primarily proposed augmenting long-tail classes by pasting them from other frames.
+While effective to some extent, we find that the augmentation strength is insufficient to counteract the heightened imbalance in our setting.
+To this end, we propose an enhanced data augmentation strategy specifically designed for class balancing.
+As for projection noise, recent approaches [58] attempt to mitigate it by interpolating pseudo pixels into unoccupied regions.
+However, this often introduces new artifacts in the range-view image.
+In contrast, tailored to the multi-range configuration of FLARES, we leverage the unused sub-clouds to increase pixel occupancy in the sampled range-view image, thereby reducing projection artifacts.
+Weighted Paste-Drop+ (WPD+) To address the problem of class imbalance, a pervasive challenge in LiDAR semantic segmentation benchmarks [3, 16], Weighted Paste-Drop (WPD) [19] technique provides a simple yet effective solution by selectively pasting pixels from rare classes while dropping pixels from abundant ones.
+Building on this approach, we introduce WPD+, a fundamentally re-engineered augmentation strategy specifically designed for FLARES.
+Unlike traditional methods that apply augmentation uniformly on range-view images [19, 25], WPD+ operates directly in 3D space, This design not only avoids repeated geometric transformations on sampled and current frames, but also harnesses multiple scenes to more effectively balance class distributions and mitigate the severe imbalance induced by point cloud splitting.
+To further enhance occurrence of underrepresented classes, particularly those corresponding to small and dynamic objects in the scene, we incorporate a curated synthetic dataset from the Carla Simulator [14] (more details are provided in Appendix A.5).
+From our empirical results, despite inherent domain gaps, this approach has led to notable accuracy gains in long-tail classes.
+Multi-Cloud Fusion (MCF) Due to the point cloud splitting in FLARES, the decrease in 2D occupancy leads to intensified projection artifacts in range images.
+To address the issue, we propose MCF, a strategy tailored to multi-range settings of FLARES.
+We formalize this process as follows.
+Let $R_{i}:\Omega\to\mathbb{R}\cup\{\varnothing\}$ denote the range image from the $i$-th sub-cloud over the pixel domain $\Omega$.
+For each pixel $x\in\Omega$, We first define the set using other range images:
+$S(x)=\{\,R_{j}(x)\mid j\neq i\text{ and }R_{j}(x)\neq\varnothing\,\}.$
+Then, for a range image $R_{i}$ over the pixel domain $\Omega$, the occupancy-filled value $\tilde{R}_{i}(x)$ is given by:
+$\tilde{R}_{i}(x)=\begin{cases}R_{i}(x),&\text{if }R_{i}(x)\neq\varnothing,\\[4.30554pt] f(S(x)),&\text{if }R_{i}(x)=\varnothing\text{ and }S(x)\neq\emptyset,\\[4.30554pt] \varnothing,&\text{otherwise.}\end{cases}$
+Here, $f(\cdot)$ is an aggregation function that combines the occupied pixel values from the remaining N1N-1 range images.
+Consequently, this method maximizes the 2D occupancy in the range image of a sub-cloud while maintaining the structural consistency of the scene.
+Despite their simplicity, both augmentation strategies are effective in enhancing segmentation performance.
+Moreover, when combined with the multi-range representations of FLARES, they significantly outperform prior methods (see Appendix B.3 for related experiments).
 3.3 Post-Processing
-After the augmented image being processed, 2D predictions from the network must be reprojected into 3D space using some post-processing technique.
-To align with the new inference framework with stacked predictions, we first propose an extension of the standard KNN method [24], termed  KNN Ensembling.
-In the post-processing phase, all sub-clouds are iteratively processed with KNN and votes are ensembled for every point from the full cloud to obtain final predictions.
-However, we found that the extension still faces the challenge in appropriately weighting contributions of nearest neighbors in 3D coordinates.
-In addition, the inference time is accumulated due to iterative process.
-To address the limitation, we propose a novel algorithm called Nearest Neighbors Range Interpolation (NNRI).
-Its pseudo-code is detailed in Algo.
+After images are processed by the network, their 2D predictions must be reprojected into 3D space using a post-processing technique.
+Conventional methods [34, 65] typically rely on hard voting from nearest neighbors to infer predictions for points.
+However, they are designed for the single range image and requires an iterative processing of all images when working on FLARES, which results in increased computational complexity.
+Furthermore, these approaches are limited in their ability to appropriately weight the contributions of each neighbor in 3D coordinates, thus leading to sub-optimal performance.
+To leverage the advantages of multi-range settings and overcome these limitations, we propose a novel algorithm called Nearest Neighbors Range Interpolation (NNRI).
+Its pseudo-code is shown in Algo.
 1.
+Define : NN sub-clouds.
+The annotation contains $C$ classes
+Input : Range images $R_{ranges}$ with size $N\times H\times W$,
+Softmax scores $I_{scores}$ with size $N\times C\times H\times W$,
+Arrays $R_{all}(p)$ with range values for all points,
+Image coordinates $(u_{all},v_{all})$ for all points,
+Kernel size $k$,
+Padding padpad,
+Cut-off factor $\alpha$,
+Mean of all range values $r_{mean}$,
+Standard Deviation of all range values $r_{std}$
+Output : Array LabelsLabels with predicted labels for all points.
+1:Unfold scores and ranges with $k\times k$ kernel:
+[…]
+[…]
+4:Extract nearest-neighbors for each point $p$:
+[…]
+[…]
+7:Compute relative depths:
+[…]
+9:Compute the cut-off value for each point $p$:
+[…]
+11:Filter the valid neighbors and compute weights:
+12: $\mathbf{N}_{valid}(n,p,k)\leftarrow\texttt{clamp}(\mathbf{N}_{rel}(n,p,k),\texttt{max}=D(p)$)
+[…]
+14:Weighted Sum for 3D Projection:
+[…]
+[…]
+17:Return LabelsLabels
+Algorithm 1 Nearest Neighbors Range Interpolation
 After applying softmax to the network output, we begin by kernelizing 2D predictions and range images using a pre-defined kernel size ($3\times 3$ in our experiments).
@@
 Finally, softmax scores of all 3D points are interpolated by the weighted sum of their nearest neighbors.
-NNRI is designed to effectively mitigate the ”many-to-one” issue inherently in range-view methods by leveraging distance-wise local neighborhood information in both 2D and 3D.
-Define : $N=N_{max}$ sub-clouds.
-The annotation contains $C$ classes
-Input : Range images $R_{ranges}$ with size $N\times H\times W$,
-Softmax scores $I_{scores}$ with size $N\times C\times H\times W$,
-Arrays $R_{all}(p)$ with range values for all points,
-Image coordinates $(u_{all},v_{all})$ for all points,
-Kernel size $k$,
-Padding $pad$,
-Cut-off factor $\alpha$,
-Mean of all range values $r_{mean}$,
-Standard Deviation of all range values $r_{std}$
-Output : Array $Labels$ with predicted labels for all points.
-1:Unfold scores and ranges with $k\times k$ kernel:
-2:    $S_{s}(n,h,w,k)\leftarrow\texttt{unfold}(I_{scores},k,pad)$
-3:    $S_{r}(n,h,w,k)\leftarrow\texttt{unfold}(R_{ranges},k,pad)$
-4:Extract nearest-neighbors for each point $p$:
-5:    $N_{s}(n,p,k)\leftarrow S_{s}(n,h,w,k)[...,u_{all},v_{all}]$
-6:    $N_{r}(n,p,k)\leftarrow S_{r}(n,h,w,k)[...,u_{all},v_{all}]$
-7:Compute relative depths:
-8:    $N_{rel}(n,p,k)\leftarrow||(N_{r}(n,p,k)-R_{all}(p)||$
-9:Compute the cut-off value for each point $p$:
-10:    $D(p)=\texttt{exp}(\frac{R(p)-r_{mean}}{r_{std}})*\alpha$
-11:Filter the valid neighbors and compute weights:
-12: $N_{valid}(n,p,k)\leftarrow\texttt{clamp}(N_{rel}(n,p,k),\texttt{max}=D(p)$)
-13:    $W(n,p,k)=1-\texttt{Normalize}(N_{valid}(n,p,k))$
-14:Weighted Sum for 3D Projection:
-15:    $Scores(p)=\sum_{i}^{k^{2}\times n}W(n,p,k)*N_{s}(n,p,k)$
-16:    $Labels=\texttt{argmax}_{c\in C}(Scores(p))$
-17:Return $Labels$
-Algorithm 1 Nearest Neighbors Range Interpolation
-3.4 Network Selection
-In order to pursue enhancement in the segmentation accuracy while possibly maintaining the high efficiency of range-view-based approaches, we revisited prior works and selected three light-weight CNN-based networks (FIDNet [48], SalsaNext [7] and CENet [5]) for integration.
-To further test the effectiveness across different network architectures, we additionally deploy RangeViT [1], a network composed of a series of Vision Transformer blocks [9], in the experimental phase.
-Original RangeViT uses a trainable KPConv-based 3D projector to get the point-wise predictions.
-We replace it with our post-processing component to achieve the full integration of our framework and train the model from scratch.
+Overall, NNRI is designed to effectively mitigate the ”many-to-one” issue inherent in range-view methods by leveraging distance-wise local neighborhood information in both 2D and 3D.
+Moreover, the new approach eliminates the need for sub-cloud concatenation, as used in STR [25], and processes all range images in parallel, which significantly reduces the inference time.
 4 Experimental Analysis
@@
 After merging similar and infrequent classes, the dataset includes 16 distinct semantic classes.
+Networks We revisited prior works and selected three light-weight CNN-based networks: FIDNet [65], SalsaNext [10] and CENet [7]) for integration.
+To further test the effectiveness across heterogeneous architectures, we additionally deploy RangeViT [2], a network composed of a series of Vision Transformer blocks [13], in the experimental phase.
+Original RangeViT uses a trainable KPConv-based 3D projector to get the point-wise predictions.
+We replace it with our post-processing component to achieve the full integration of our framework and train the model from scratch.
 Implementation Details Prior works experimented mostly with the resolution of $64\times 2048$ for SemanticKITTI [7, 34, 10], and $32\times 960$ [25] or $32\times 2048$ [2] for nuScenes.
-In contrast, we reduce the azimuth resolution while increasing the projection rate in FLARES mode: resolutions of $64\times 512$ for SemanticKITTI and $32\times 480$ for nuScenes are fixed for the input and the full point cloud is split into up to 3 and 2 ($N_{max}$) sub-clouds during training and inference.
-We directly use custom configurations of prior works [48, 5, 7, 1] to train the networks in FLARES mode.
+In contrast, we reduce the azimuth resolution while increasing the projection rate in FLARES mode: resolutions of $64\times 512$ for SemanticKITTI and $32\times 480$ for nuScenes are fixed for the input and the full point cloud is split into up 3 and 2 sub-clouds during training and inference, respectively.
+FLARES uses the same loss configurations of prior works [65, 7, 10, 2], including a cross-entorpy loss (focal loss for RangeViT) and Lovász–Softmax loss [4].
 For training the selected models (excluding RangeViT) on the nuScenes dataset, we standardize the hyperparameter set since no default configurations are provided.
@@
 All models are trained on four NVIDIA GeForce GTX 1080Ti in distributed mode.
-Evaluation Metrics Following prior works, we assess the performance using Intersection-over-Union (IoU) $\text{IoU}_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ and Accuracy (Acc) $\text{Acc}_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}}$ for each class $i$, and calculate the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) across all classes.
-$TP_{i}$, $FP_{i}$, and $FN_{i}$ represent the true positives, false positives, and false negatives for class $i$, respectively.
 4.2 Comparative Study
@@
 1, all four networks see significant improvements: SalsaNext has mIoU gains of 5.3% on SemanticKITTI and 2.3% on nuScenes, FIDNet improves by 7.9% and 3.9%, and CENet by 3.3% and 3.1%.
-RangeViT, as a ViT-based network, also exhibits huge enhancement in performance, confirming FLARES’s generalizability across different architectures.
+RangeViT, as a ViT-based network, also exhibits huge enhancement in performance, confirming FLARES’s generalization across different architectures.
 This improvement is especially prominent for smaller, dynamic, and under-represented classes such as truck, motorcycle, bicycle, pedestrian and bicyclist.
-In Fig.
-4, we present a visual comparison.
 Notably, with the support of FLARES, the network demonstrates improved accuracy in segmenting foreground objects.
+For instance, as shown in Fig.
+4, our method’s predictions exhibit significantly higher correctness compared to the baseline.
 An exception arises with the motorcyclist class in SemanticKITTI, where IoU scores decrease compared to the baseline.
@@
 We regard this as a corner case when testing on an class-imbalanced dataset.
-As a future work to resolve the issue, we aim to explore 3D reconstruction techniques to generate real-world-like pseudo LiDAR point clouds for augmentation [4, 23].
+As a future work to resolve the issue, we aim to explore 3D reconstruction techniques to generate real-world-like pseudo LiDAR point clouds for enhanced augmentation [5, 33].
 In Tab.
-2, we further compare FLARES-boosted networks with other state-of-the-arts from various modalities.
-Given that methods we select preserve relatively fewer number of parameters, FLARES helps to improve the segmentation accuracy, being comparable to other point- or voxel-based approaches that deploy much larger and deeper networks.
-In addition, they outperform others significantly in latency, achieving superior trade-offs in accuracy and efficiency.
+2, we compare the performance of networks boosted by FLARES with other state-of-the-art approaches across various modalities.
+Despite using relatively fewer parameters, FLARES-enhanced models achieve segmentation accuracy comparable to other point- or voxel-based methods that deploy much larger and deeper networks.
+Moreover, our approach significantly outperforms these alternatives in terms of latency, achieving an excellent trade-off between accuracy and efficiency.
  SemanticKITTI test set
@@
 SalsaNext [10]     $59.5$     $91.9$ $48.3$ $38.6$ $38.9$ $31.9$ $60.2$ $59.0$ $\mathbf{19.4}$ $91.7$ $63.7$ $75.8$ $\mathbf{29.1}$ $90.2$ $64.2$ $81.8$ $63.6$ $\mathbf{66.5}$ $54.3$ $62.1$
-[…]
-[…]
+ΥSalsaNextΥ     $63.3$     $94.7$ $52.9$ $55.7$ $57.3$ $50.2$ $65.5$ $70.9$ $13.0$ $92.6$ $\mathbf{69.0}$ $77.7$ $20.5$ $90.4$ $65.8$ $80.8$ $65.0$ $63.4$ $55.4$ $62.4$
+ΥSalsaNextΥ     $\mathbf{64.8}$     $\mathbf{95.1}$ $\mathbf{55.5}$ $\mathbf{56.5}$ $\mathbf{60.1}$ $\mathbf{53.7}$ $\mathbf{69.6}$ $\mathbf{74.1}$ $11.4$ $\mathbf{93.0}$ $68.9$ $\mathbf{78.9}$ $20.4$ $\mathbf{91.1}$ $\mathbf{67.6}$ $\mathbf{82.0}$ $\mathbf{66.7}$ $65.0$ $\mathbf{58.1}$ $\mathbf{64.1}$
 FIDNet [65]     $51.3$     $90.4$ $28.6$ $30.9$ $34.3$ $27.0$ $43.9$ $48.9$ $16.8$ $90.1$ $58.7$ $71.4$ $19.9$ $84.2$ $51.2$ $78.2$ $51.9$ $64.5$ $32.7$ $50.3$
 FIDNet     $59.5$     $93.9$ $54.7$ $48.9$ $27.6$ $23.9$ $62.3$ $59.8$ $\mathbf{23.7}$ $90.6$ $59.1$ $75.8$ $26.7$ $88.9$ $60.5$ $84.5$ $64.4$ $69.0$ $53.3$ $62.8$
-[…]
-[…]
+ΥFIDNetΥ     $65.1$     $95.3$ $51.0$ $57.0$ $54.8$ $58.1$ $68.1$ $68.9$ $14.4$ $92.3$ $68.3$ $78.0$ $32.3$ $91.6$ $67.6$ $83.7$ $66.6$ $68.8$ $55.1$ $64.8$
+ΥFIDNetΥ     $\mathbf{67.4}$     $\mathbf{95.8}$ $\mathbf{56.7}$ $\mathbf{60.7}$ $\mathbf{58.1}$ $\mathbf{60.3}$ $\mathbf{72.5}$ $\mathbf{72.9}$ $15.8$ $\mathbf{93.2}$ $\mathbf{69.2}$ $\mathbf{79.9}$ $\mathbf{34.2}$ $\mathbf{91.9}$ $\mathbf{69.0}$ $\mathbf{84.6}$ $\mathbf{68.7}$ $\mathbf{70.3}$ $\mathbf{59.9}$ $\mathbf{66.9}$
 ⋄†CENet [7]     $60.7$     $92.1$ $45.4$ $42.9$ $43.9$ $46.8$ $56.4$ $63.8$ $29.7$ $91.3$ $66.0$ $75.3$ $31.1$ $88.9$ $60.4$ $81.9$ $60.5$ $67.6$ $49.5$ $59.1$
 CENet     $64.7$     $91.9$ $58.6$ $50.3$ $40.6$ $42.3$ $68.9$ $65.9$ $\mathbf{43.5}$ $90.3$ $60.9$ $75.1$ $31.5$ $91.0$ $66.2$ $84.5$ $69.7$ $\mathbf{70.0}$ $61.5$ $67.6$
-[…]
-[…]
+ΥCENetΥ     $66.6$     $95.6$ $58.5$ $61.6$ $51.7$ $50.2$ $74.5$ $72.4$ $23.2$ $91.4$ $69.6$ $77.1$ $31.7$ $91.1$ $66.6$ $83.8$ $69.9$ $68.3$ $60.3$ $68.7$
+ΥCENetΥ     $\mathbf{68.0}$     $\mathbf{95.9}$ $\mathbf{61.1}$ $\mathbf{62.1}$ $\mathbf{57.2}$ $\mathbf{59.0}$ $\mathbf{77.2}$ $\mathbf{74.2}$ $12.2$ $\mathbf{92.2}$ $\mathbf{69.9}$ $\mathbf{78.7}$ $\mathbf{32.9}$ $\mathbf{91.8}$ $\mathbf{68.8}$ $\mathbf{84.7}$ $\mathbf{71.3}$ $69.9$ $\mathbf{62.9}$ $\mathbf{70.3}$
 RangeViT [2]     $64.0$     $95.4$ $55.8$ $43.5$ $29.8$ $42.1$ $63.9$ $58.2$ $\mathbf{38.1}$ $\mathbf{93.1}$ $70.2$ $\mathbf{80.0}$ $32.5$ $\mathbf{92.0}$ $\mathbf{69.0}$ $\mathbf{85.3}$ $\mathbf{70.6}$ $\mathbf{71.2}$ $\mathbf{60.8}$ $64.7$
-[…]
+ΥRangeViTΥ     $\mathbf{66.1}$     $\mathbf{95.6}$ $\mathbf{56.3}$ $\mathbf{60.5}$ $\mathbf{52.4}$ $\mathbf{57.1}$ $\mathbf{72.0}$ $\mathbf{69.7}$ $16.0$ $91.6$ $\mathbf{71.1}$ 77.3{77.3} $\mathbf{32.7}$ $91.4$ $67.4$ $83.1$ $68.0$ $68.1$ $58.0$ $\mathbf{67.5}$
  nuScenes val set
@@
 SalsaNext     $72.2$     $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $\mathbf{63.1}$ $61.3$ $76.5$ $96.0$ $\mathbf{70.8}$ $71.2$ $71.5$ $86.7$ $84.4$
-[…]
+ΥSalsaNextΥ     $\mathbf{74.5}$     $\mathbf{75.0}$ $\mathbf{34.6}$ $\mathbf{90.4}$ $\mathbf{90.0}$ $\mathbf{43.8}$ $\mathbf{79.4}$ $\mathbf{72.9}$ $58.8$ $\mathbf{65.8}$ $\mathbf{79.9}$ $\mathbf{96.5}$ $70.1$ $\mathbf{74.0}$ $\mathbf{73.9}$ $\mathbf{87.6}$ $\mathbf{85.6}$
 FIDNet     $72.7$     $73.0$ $36.0$ $87.8$ $86.0$ $45.6$ $74.1$ $73.9$ $62.5$ $67.1$ $77.7$ $94.3$ $69.8$ $72.2$ $72.1$ $86.1$ $84.5$
-[…]
+ΥFIDNetΥ     $\mathbf{76.6}$     $\mathbf{77.6}$ $\mathbf{43.5}$ $\mathbf{92.9}$ $\mathbf{88.1}$ $\mathbf{56.5}$ $\mathbf{79.5}$ $\mathbf{77.7}$ $\mathbf{65.3}$ $\mathbf{67.0}$ $\mathbf{83.1}$ $\mathbf{96.6}$ $\mathbf{72.8}$ $\mathbf{75.0}$ $\mathbf{74.5}$ $\mathbf{88.5}$ $\mathbf{86.8}$
 CENet     $73.7$     $73.6$ $32.9$ $92.7$ $87.1$ $\mathbf{53.5}$ $76.1$ $69.0$ $58.7$ $66.8$ $81.6$ $95.6$ $71.1$ $73.7$ $73.2$ $87.5$ $85.7$
-[…]
+ΥCENetΥ     $\mathbf{76.8}$     $\mathbf{76.7}$ $\mathbf{45.2}$ $\mathbf{93.5}$ $\mathbf{90.3}$ $49.6$ $\mathbf{83.1}$ $\mathbf{78.1}$ $\mathbf{66.4}$ $\mathbf{69.0}$ $\mathbf{82.5}$ $\mathbf{96.6}$ $\mathbf{73.9}$ $\mathbf{75.1}$ $\mathbf{74.6}$ $\mathbf{88.3}$ $\mathbf{86.3}$
 RangeViT     $75.2$     $75.5$ $\mathbf{40.7}$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $\mathbf{66.3}$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $\mathbf{89.9}$ $\mathbf{87.2}$
-[…]
+ΥRangeViTΥ     $\mathbf{77.0}$     $\mathbf{76.7}$ $39.2$ $\mathbf{93.0}$ $\mathbf{92.0}$ $\mathbf{55.2}$ $\mathbf{81.6}$ $\mathbf{77.2}$ $64.9$ $\mathbf{70.9}$ $\mathbf{84.1}$ $\mathbf{96.8}$ $\mathbf{74.1}$ $\mathbf{75.6}$ $\mathbf{75.1}$ $88.6$ $86.7$
 Table 1: Comparisons of state-of-the-art LiDAR semantic segmentation methods on the test set of SemanticKITTI [3] and val set of nuScenes [16] in standard and  FLARES mode.
@@
 […]
-[…]
+ΥSalsaNextΥ    6.7M     29     Range     74.5     64.8     64.8
 PolarNet [63] [’20]    13.6M     71     Polar     71.0     54.9     57.5
@@
 Tornado-Net [18] [’20]    -     -     Multiple     -     64.5     63.1
-[…]
+ΥFIDNetΥ    6.1M     26     Range     76.6     65.6     67.4
 Cylinder3D [66] [’21]    56.3M     170     Voxel     76.1     67.8     65.9
@@
 Lite-HDSeg [41] [’21]    -     50     Range     -     64.4     63.8
-[…]
+ΥCENetΥ    6.8M     24     Range     76.8     67.5     68.0
 Meta-RSeg [48] [’22]    6.8M     46     Range     -     60.3     61.0
@@
 GFNet [39] [’22]    -     100     Multiple     76.8     63.2     65.4
+WaffleIron [37] [’23]    6.8M     143     Point     79.1     68.0     70.8
+SphereFormer [26] [’23]    32.3M     165     Multi     78.4     67.8     74.8
+PTv3 [53] [’24]    46.2M     67     Point     80.4     70.8     74.2
 […]
 4.3 Ablation Study
-To perform the ablation study, we test with CENet [5] on val set of SemanticKITTI [2].
-Training Schema As shown in Tab.3, we conduct a comprehensive evaluation to examine the trade-off between accuracy and efficiency under different training configurations.
-In the standard mode, a high-resolution range image is used as the input to reproduce baseline results.
-While training in the STR paradigm[18], there is a slight performance drop, the memory consumption is significantly reduced by partitioning the full point cloud through azimuth-wise grouping.
-The limited increase in latency is due to the fact that STR has to additionally unite the batched prediction along azimuth resolution and the post-processing is still performed on the high-resolution image.
-In contrast, FLARES achieves remarkable improvements, showing increases of 2.7% in mIoU, 2.1% in mAcc, and 45% acceleration in inference.
-Similar to STR, FLARES reduces memory consumption compared to the standard mode, further optimizing the efficiency of range-view semantic segmentation.
-Method Input resolution mIoU mAcc Lat.
-Standard $1\times 64\times 2048$ 64.8 77.2 44 ms
-STR [18] $1(5)\times 64\times 480$ 64.3 76.8 41 ms
-FLARES $1(3)\times 64\times 512$ 67.5 79.3 24 ms
-Table 3: Performance with varying training schemes is evaluated on val set of SemanticKITTI [2].
-Input resolution is formatted in $B\times H\times W$ (batch size, image width and height).
-For STR [18], we use the configuration yielding the best performance on CENet [5]: a full resolution of $64\times 1920$ split into four $64\times 480$ sub-images.
-For the fair comparison, all models are integrated with proposed components and trained from scratch.
-Data Augmentation As presented in Fig.
-5(a), various data augmentation methods can enhance the segmentation performance by large margins and our WPD+ performs the best among them, which increases baseline mIoU by 5.7% and mAcc by 2.9%.
-Mix3D [25] increases the contextual information per frame by fusing one scene to another, however, the model can still suffer from class imbalance.
-RangeAug [18] consists of 4 different range-wise operations to enrich the semantic and structural cues.
-From their experiments, it demonstrated that this augmentation technique is especially effective for attention-based networks [33, 9], which possess significantly higher model capacities and are more reliant on data diversity for optimal performance.
-Similar to their experimental results, applying this method to lightweight CNN-based networks has shown limited success in improvement.
-(a) WPD+
-(b) #Frames sampled in WPD+
-(c) Inclusion of synthetic dataset
-(d) MCF
-Figure 5: a) Initialization of training with standard geometric data augmentation (GDA) and benchmark several state-of-the-art 3D augmentation techniques, including Mix3D [25], RangeAug [18], and original WPD [13].
-b) Different number of sampled frames for best performance c) A comparative plot showing the IoU scores of top-rare classes in scenarios both with and without the inclusion of the synthetic dataset.
-As reference, the class frequencies in the val set are provided.
-d) The models are trained using two different input configurations: either a single range image derived from the full cloud (FC) or a range image generated from a sub-cloud (SC), and inferred in FLARES mode.
-Note that for all trained models in a) - d), we leverage KNN Ensembling during post-processing phase.
-WPD+ includes two tunable parameters: the number of sampled frames from the original dataset and the use of the synthetic dataset.
-To find out the best parameter set, we conduct two additional experiments.
+To perform the ablation study, we test with CENet [7] on val set of SemanticKITTI [3] dataset.
+STR [25]     FLARES     WPD+     MCF     NNRI     mIoU     Lat.
+                    63.1     44 ms
+✓                     62.6     41 ms
+    ✓                 64.2     46 ms
+    ✓     ✓             65.5     -
+    ✓     ✓     ✓         65.9     -
+    ✓     ✓     ✓     ✓     67.5     24 ms
+Table 3: Full ablation study on SemanticKITTI [3] dataset.
+The first row denotes the baseline results trained in standard mode with high-resolution range images ($64\times 2048$).
+For STR [25], we split the full range image into 4 sub-images, each has the resolution of $64\times 512$.
+For all experiments without NNRI, we employ the conventional KNN [34] post-processing.
+Component Design In Tab.
+3, we evaluate the contributions of each component in FLARES.
+Starting with the baseline results, we observe that applying STR [25] slightly reduces mIoU with the limited improvement in latency.
+Incorporating FLARES into the framework then yields a significant boost in performance compared to the baseline.
+However, since KNN post-processing must be iteratively applied to multiple range images, the inference speed becomes lower.
+Next, our two data augmentation methods effectively addressing the exacerbated class imbalance and amplified noise introduced by FLARES.
+Consequently, this leads to a further increase of 1.7% in mIoU.
+Finally, replacing the standard post-processing with NNRI not only improves the overall performance but also achieves a remarkable speed-up in inference due to its ability to run parallel computations across multiple range images.
+Essentially, compared to the baseline performance, we achieve 4.4% mIoU gain and around 45% speed-up through the new framework.
+Figure 5: a) Ablation on the number of sampled frames for WPD+.
+b) Comparative plot of IoU scores for top rare classes with and without the synthetic dataset, alongside the class frequency distribution in the validation set.
+For all inferences, KNN [34] is used as the post-processing approach.
+WPD+ To address the challenge of class imbalance within FLARES, WPD+ is incorporated as a data augmentation step.
+It includes two tunable parameters: the number of sampled frames from the original dataset and the inclusion of the synthetic dataset.
+To determine the optimal configuration, we conducted two ablation experiments.
 Fig.
-5(b) shows that sampling 6 frames results in the optimal performance, while increasing the number of frames beyond this point leads to performance degradation.
-This is likely due to the model’s limited scalability, similar to the effects observed in case of RangeAug[18], in addition, pasting too many pixels belonging to specific classes will again corrupt the semantic balance.
+5(a) shows that sampling 6 frames results in the optimal performance, while increasing the number of frames beyond this point leads to performance degradation.
+Fusing too many frames can introduce noise and redundant information that may overwhelm the model, for instance, excessive fusion might result in overlapping objects and distorted boundaries when too many new objects are pasted into the scene.
 Furthermore, Fig.
-5(c) illustrates that the synthetic dataset plays a key-role in refining semantic prediction of top-rare classes.
-Noteworthily, using the synthetic dataset is efficient and practical because this allows us to customize sensor configurations to align with the target dataset and to define specific objects within the scene for downstream applications without any labor cost.
-In the next stage, we study how unoccupied pixels in the range image can affect the performance.
-Firstly, we use the full point cloud for training and sub-clouds for inference, as the decreasing azimuth resolution can already resolve the problem of low 2D occupancy (the first column in Fig.
-5(d)).
-However, empirical results reveal a drop in 2D accuracy when inferring from sub-clouds.
-This is possibly because of the domain shift caused by the occupancy difference in range images.
-In the next trial, we use the sub-cloud as training input instead, but this directly limits the performance due to the occupancy reduction.
-To tackle this compounded challenge brought by FLARES, we introduce Multi-Cloud Fusion, an additional data augmentation step which fuses multiple sub-clouds through occupancy padding during projection phase.
-As shown in Fig.
-5(d), using MCF exhibits the highest performance in the series, achieving 1% increase in IoU over the model trained on full cloud.
-As an alternative, we apply RangeIP [44], an interpolation-based augmentation technique, to enhance 2D occupancy in the range image, but it results in slight worse accuracy compared to the baseline.
-(a) Ablation on Post-Processing
-(b) Inference Time [ms]
-(c) Ablation on Input Resolution
-Figure 6: a) - b): Various post-processing techniques are applied to the same trained model.
-KNN [24] and NLA [48] are used on the single range image generated from the full point cloud, whereas KNN Ensembling operates on multiple range images derived from sub-clouds.
-For NNRI, we assess it using both full cloud and sub-clouds.
-c): Evaluation on different input resolutions and corresponding number of sub-clouds ($N_{max}$).
-Post-Processing As introduced in Sec.
-3.3, a simple LiDAR model is built to approximate the distance-density function of 3D points and compute cut-off values for valid neighbors extraction.
-To verify the necessity of the step, some qualitative results are provided in Fig.
-7.
-As can be seen, the adaptive cut-off values refine semantic predictions by better accommodating objects with varying density scales.
+5(b) illustrates that the synthetic dataset plays a key-role in refining semantic prediction of top-rare classes.
+Notably, using the synthetic dataset is both efficient and practical, as it allows us to customize sensor configurations to align with the target dataset and define specific objects within the scene for downstream applications without incurring any labor cost.
+Figure 6: Ablation on various post-processing approaches: KNN [34], NLA [65], and NNRI are first compared in a single-range setting.
+*NNRI denotes the multi-range variant of our post-processing method.
+For fair evaluation, we also extend KNN to operate on multiple range images (*KNN).
+Dotted lines indicate the metrics for 2D predictions.
+Post-Processing We explore the performance of various post-processing techniques on both efficacy and efficiency in Fig.
+6.
+Regarding conventional KNN [34] as the baseline, NLA [65] demonstrates similar performance in both accuracy and latency.
+In contrast, we deploy our approach (NNRI) in the standard setting as well and observe a significant improvement: inference time is cut nearly 16% compared to KNN, while mAcc and mIoU increase by 2.8% and 2%, respectively.
+Unlike KNN, NNRI avoids the computational cost of Gaussian kernel calculations for distance weighting and directly performs nearest neighbor searches on the range image instead of in 3D space, which further reduces computational overhead.
+NNRI interpolates class-wise scores based on relative depths rather than directly voting on hard labels, relying more on weighted information from nearest neighbors, which is the major reason why it outperform other post-processing approaches.
+Next, switching to FLARES mode, we first implement an extension of KNN, which iteratively gathers votes from all points in each sub-cloud and aggregates them for the final prediction.
+While this extension improves the accuracy, it comes at approximately doubled latency cost.
+Conversely, when NNRI is adapted to all sub-clouds, it consistently provides notable improvements in both efficacy and efficiency.
+As a reference, we included evaluation scores on 2D predictions, showing that FLARES with NNRI significantly narrows the accuracy gap between 2D and 3D predictions.
+This suggests that our approach effectively mitigates the ”many-to-one” problem and offers substantial gains in segmentation performance.
 (a) Scene
@@
 Similarly, in the bottom image, half of the points that belong to Building are incorrectly predicted as Vegetation.
-Next, we explore the impact of various post-processing techniques on segmentation performance in Fig.
-6(a) and 6(b).
-Regarding conventional KNN [24] as the baseline, NLA [48] demonstrates similar performance in both accuracy and latency.
-In contrast, we deploy our approach (NNRI) in the standard mode as well and observe a significant improvement: inference time is cut nearly 16% compared to KNN, while mAcc and mIoU increase by 2.8% and 2%, respectively.
-Unlike KNN, NNRI avoids the computational cost of Gaussian kernel calculations for distance weighting and directly performs nearest neighbor searches on the range image instead of in 3D space, further reducing computational overhead.
-NNRI interpolates class-wise scores based on relative depths rather than directly voting on hard labels, relying more on weighted information from nearest neighbors, which is the major reason why it outperform other post-processing approaches.
-Switching to FLARES mode, we first extend the standard KNN approach to KNN Ensembling, which iteratively gathers votes from all points in each sub-cloud and aggregates them for the final prediction.
-While this extension improves the accuracy, it comes at approximately doubled latency cost.
-Conversely, when NNRI is adapted, it consistently provides notable improvements in both efficacy and efficiency.
-As a reference, we included evaluation scores on 2D predictions (-- dashed lines in Fig.
-6(a)), showing that FLARES with NNRI significantly narrows the accuracy gap between 2D and 3D predictions.
-This suggests that our approach effectively mitigates the ”many-to-one” problem, offering substantial gains in segmentation performance.
-Input Resolution To further explore the optimal results in FLARES mode, we test various input resolutions, as shown in Fig.
-6(c).
-From the experimental results, we found that resolutions of $512$ and $640$ deliver the best mIoU scores.
-Increasing the azimuth resolution beyond this point causes a slight performance drop.
-Nevertheless, all tested configurations outperform the baseline (the first row in Tab.3), demonstrating the superb effectiveness of our approach on range-view LiDAR semantic segmentation.
+Furthermore, we optimize the implementation of NNRI by leveraging an adaptive cut-off value to filter valid nearest neighbors.
+This parameter is derived by approximating the internal LiDAR geometry and the distance-density relationship of 3D points.
+To illustrate its impact, qualitative results are provided in Fig.
+7.
+As shown, the adaptive cut-off refines semantic predictions by better accommodating objects with varying density scales.
+Figure 8: (a) Ablation on number of sub-clouds (NN) with fixed image resolution of $64\times 512$. (b) Ablation on image width $W$ (NN is fine-tuned for each configuration, respectively).
+Input Configuration In Fig.
+8(a), we investigate in the effect on the number of splitting.
+It shows that increasing the number of sub-clouds beyond 4 degrades performance due to insufficient occupancy and projection artifacts.
+Conversely, fewer sub-clouds speeds up the inference but compromises accuracy.
+Overall, our choice of 3 sub-clouds optimally balances efficiency and accuracy.
+In Fig 8(b), we further test various azimuth resolution of input images.
+The experimental results indicate that width of $512$ delivers the best mIoU scores.
+Increasing or decreasing the azimuth resolution beyond this point causes a slight performance drop.
 5 Conclusion
-In this work, we introduced FLARES, an optimized training and inference schema designed for seamless integration into any range-view-based network.
-To enhance the scalability, we further developed two data augmentation techniques for class rebalance and occupancy increase.
-Additionally, we proposed a novel unsupervised post-processing method to effectively address the ”many-to-one” issue, minimizing the 2D-to-3D performance gap.
-Our approach has substantiated significant improvements in accuracy and efficiency over baselines across various network architectures on two widely used LiDAR benchmarks.
-Despite the limitation in some corner case of class imbalance in the dataset, our approach shows its overall versatility and effectiveness in advancing LiDAR semantic segmentation.
+In this work, we introduce FLARES, an optimized multi-range training paradigm designed for range-view LiDAR semantic segmentation.
+Our framework seamlessly integrates into any range-view-based network.
+We also develop two data augmentation techniques to address the challenges of exacerbated class imbalance and amplified projection noise within FLARES.
+Additionally, we propose a novel post-processing method tailored to multi-range settings to effectively tackle the ”many-to-one” issue.
+Our approach yields significant improvements in both accuracy and efficiency over baselines across various network architectures on two widely used LiDAR benchmarks.
+The limitations of our current work are twofold: (1) although our data augmentation is powerful, it remains limited in corner cases of extremely low occurrence of certain classes, motivating further exploration of solutions to class imbalance; and (2) the fixed nature of the splitting step may result in the loss of critical information.
+In future work, we aim to address these limitations and develop a more robust framework that can be generalized to more challenging scenarios, such as adverse weather conditions or limited annotations.
+Supplementary Material
+Appendix A More Implementation details
+In this section, we provide additional details regarding to the overall workflow, proposed components and configurations during training and inference.
+Evaluation Metrics Following prior works, we assess the performance using Intersection-over-Union (IoU) and Accuracy (Acc) for each class, and calculate the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) across all classes.
+A.1 Detailed Overview
+Each step in the training and testing process is detailed in Fig.
+9.
+To maximize the efficiency, the sequence of pre-processing and data augmentation steps is fixed.
+WPD+ is applied first, before Geometric Data Augmentation (GDA), to avoid performing the same geometric transformation on all sampled frames.
+Once these two point-level steps are completed, the augmented point cloud is downsampled and divided into multiple sub-clouds, on which projection is performed.
+Multi-Cloud Fusion (MCF) is applied last, conducting pixel-wise fusion on the range images from each sub-cloud.
+The same steps are repeated for ground-truth labels to obtain their 2D representation.
+The final input and ground-truth label are then used to train the 2D semantic segmentation network.
+During testing, only the pre-processing steps are included, and all range images are fed into the network.
+Ultimately, we apply NNRI to upsample 2D outputs to gather 3D predictions in an unsupervised manner.
+Figure 9: Detailed workflow overview: During training, the raw point cloud is pre-processed and augmented to generate the range image.
+A 2D segmentation network is then trained to predict 2D semantic labels.
+In testing, all augmentation steps are removed, and range images are stacked as a batch.
+The 2D predictions are then gathered and processed by NNRI, our proposed post-processing component, to map them into 3D space.
+A.2 Geometric Data Augmentation
+The standard augmentation is utilized commonly in the previous works [25, 7, 2]. Hence, we follow the default configurations and include the following geometric transformations in our pipeline:
+  • Random Flipping The point cloud is randomly flipped along the $X$-axis.
+  • Random Translation The point cloud is randomly jittered in translation. The jittering values range from -5m to 5m, -3m to 3m, and -1m to 0m for $X$-axis, $Y$-axis, and $Z$-axis, respectively.
+  • Random Rotation The point cloud is randomly and globally rotated along yaw, pitch and roll direction. The ranges are set at [-5, 5] in degrees for all axes.
+  • Probability the value of randomness is set at 0.5 for all augmentation components.
+A.3 Runtime Measurement
+In Table 2, we report the inference time of various methods.
+All measurements are conducted using a single NVIDIA GTX 1080Ti GPU with a fixed batch size of 1.
+To eliminate the impact of initialization overhead, we discard the first 100 iterations and compute the average runtime over the subsequent 100 iterations.
+For methods that could not be measured directly due to GPU limitations, we report the inference times provided in their original papers, most of which were obtained using more powerful GPUs.
+Notably, all our reported times include data loading, pre-processing, and post-processing steps to ensure a fair comparison.
+A.4 Weighted Paste Drop+
+We directly follow the configuration of Weighted Paste Drop (WPD) introduced by Gu et al.
+[19]: we first compute the normalized weights from the class-wise frequencies and set the threshold of classifying long-tail and non-long-tail classes at 0.1.
+More specifically, we paste the points of classes with weights greater than 0.1 and drop them vice versa.
+The normalized weight of each class is directly utilized as the probability of pasting and dropping.
+For reference, we provide the relevant statistics on SemanticKITTI [^2] https://github.com/PRBonn/semantic-kitti-api [3] and nuScenes [^3] https://www.nuscenes.org/nuscenes [16] in Fig.
+11.
+Figure 10: An example of pasting points of long-tail classes from synthetic data onto real-world data is shown, with the fused point cloud and range image visualized in the bottom right.
+Since WPD+ does not account for spatial differences in scene-related contexts, some unrealistic scenarios may arise, such as a motorcyclist riding on a non-drivable surface.
+Nevertheless, these augmentation steps enhance the semantic richness of the scene.
+Given that long-tail classes typically correspond to small and dynamic objects, the negative impact of domain differences remains minimal in this case.
+(a) SemanticKITTI [3]
+(b) nuScenes [16]
+Figure 11: Statistical results of class-wise normalized weights on two datasets.
+When applying WPD+ to the point clouds in the SemanticKITTI dataset, we observe that the resulting range images sometimes contain artifacts outside of the contextual distributions.
+As shown in Fig.
+12, points at farther distances are left unlabeled in the dataset, as they are less relevant for most perceptual tasks.
+However, these points can introduce additional noise into the projection when non-long-tail foreground points are excluded.
+Such outliers can adversely impact training stability, so we further clean the projected range images by filtering out unlabeled representations when training the model with SemanticKITTI dataset.
+Figure 12: An example of dropping points of non-long-tail classes in a point cloud from SemanticKITTI [3] dataset: since the dataset is only annotated within the range of 50 meters, points that are out of this range can result in some artifacts in the projection after points of foreground objects are dropped in the scene.
+Hence, we insert an additional cleaning step after projection for the dataset.
+A.5 Synthetic Dataset
+In addition to copy-pasting points of long-tail classes from the same dataset, we use Carla Simulator [14] to further enhance class balance. This subsection details the process of data collection.
+Since WPD+ pastes points without spatial transformation and does not account for sensor difference, it is essential that the simulator precisely matches the sensor configuration of the target dataset to prevent introducing out-of-context points during augmentation.
+We therefore configure the simulator separately for each dataset, based on their respective sensor specifications, as detailed in Tab.
+4.
+Next, we set up the scene-related contexts.
+Based on statistics shown in Fig.
+11, we initialize the scene with 20 dynamic actors randomly selected from vehicle categories busses, motorcycles and bicycles in the Carla Simulator, along with 30 pedestrians with random movement.
+Town10 serves as the static background map for the scene, and we record 2000 frames each for SemanticKITTI and nuScenes.
+Dataset     V.FoV [deg]     H.FoV [deg]     Range [m]
+Sem.KITTI     $25\sim 3$     $180\sim 180$     $2\sim 50$
+nuScenes     $30\sim 10$     $180\sim 180$     $2\sim 80$
+Table 4: LiDAR specifications of two datasets.
+After collecting the dataset, we manually define a class mapping scheme between the real and synthetic datasets.
+The complete mapping function is detailed in Tab.
+5.
+Generally, positive correspondences can be identified; however, an exception exists where the rider class in the synthetic dataset represents both motorcyclist and bicyclist.
+To make more effective use of the synthetic dataset, we split and reassign the labels for this class.
+Since rider always appears alongside motorcycle and bicycle in the scene, we assign new labels by finding the nearest neighbor in 3D space for each rider point and categorizing it in a binary manner.
+We provide an example of the pasting process with the synthetic data in Fig.
+10 for the comprehensive understanding.
+Sem.KITTI Carla
+bicycle bicycle
+motorcycle motorcycle
+truck truck
+other vehicle bus, train
+person pedestrian
+bicyclist rider*
+motorcyclist rider*
+other ground ground
+trunk -
+pole pole
+traffic sign traffic sign
+nuScenes Carla
+barrier -
+bicycle bicycle
+bus bus
+const. vehicle -
+motorcycle motorcycle
+pedestrian pedestrian
+traffic cone -
+trailer -
+other flat ground
+Table 5: Class mapping dictionaries: we map only the long-tail classes and omit mappings where there is ambiguity in finding corresponding class names.
+A.6 Multi-Cloud Fusion
+To better illustrate the impact of Multi-Cloud Fusion (MCF) during data augmentation, we provide an example in Fig.
+13.
+Notably, range images with high azimuth resolution tend to have a greater number of empty pixels, which are distributed randomly and can distort object geometry within the scene.
+Reducing the azimuth resolution produces a more complete range image; however, occupancy declines again if only the sub-cloud is available for projection.
+MCF addresses this issue by filling unoccupied pixels, revisiting data from other sub-clouds at the same 2D positions.
+In this way, MCF effectively mitigates the problem while preserving the underlying geometry of the scene.
+(a) FC, $64\times 2048$
+(b) FC, $64\times 512$
+(c) SC, $64\times 512$
+(d) SC, $64\times 512$ (MCF)
+Figure 13: Visualization of cropped range images with varying resolutions and configurations of point clouds.
+Here, FC denotes the full cloud, and SC indicates the sub-cloud.
+The final image is augmented with Multi-Cloud Fusion (MCF).
+A.7 Multi-range KNN
+The KNN-based post-processing approach [34] has been widely used in prior works.
+To apply this method with the FLARES setup, we extend it to handle multiple range images, allowing us to infer 3D semantic labels for a single full point cloud.
+The pseudo code is provided in Algo.
+2.
+Briefly, we gather KNN votes for all points from each image iteratively, accumulating them across images.
+Each point is then assigned the class with the highest vote count.
+Define : $N=N_{max}$ sub-clouds.
+The annotation contains $C$ classes
+Input : Range images $R_{ranges}$ with size $N\times H\times W$,
+Label images $L_{labels}$ with size $N\times H\times W$,
+Arrays $R_{all}(p)$ with range values for all points,
+Image coordinates $(u_{all},v_{all})$ for all points,
+Output : Array LabelsLabels with predicted labels for all points.
+[…]
+foreach $i$ in $1:N$ do
+  $R_{range}=R_{ranges}(i)$
+  $L_{label}=L_{labels}(i)$
+  $Vote(p)\mathrel{+}=\mathbf{KNN}(R_{range},L_{label},R_{all},(u_{all},v_{all}))$
+end foreach
+$Labels=\texttt{argmax}_{c\in C}(Votes(p))$
+Algorithm 2 Multi-range KNN
+A.8 Pre-training
+In our main experiments, we test the effectiveness of FLARES using four different networks: SalsaNext [10], FIDNet [65], CENet [7], and RangeViT [2].
+For specific network designs, we refer readers to the respective original papers.
+Notably, RangeViT [2] demonstrated the advantages of using pretrained models on image datasets [42].
+Following this approach, we experiment with initializing network weights pre-trained on Cityscapes [9].
+While we apply this pre-training strategy to the other three convolution-based networks, it yields only minor improvements compared to random initialization.
+We hypothesize that Convolutional Neural Networks (CNNs) are less effective than Vision Transformers (ViTs) in terms of transferability and scalability, as ViTs are better equipped to learn long-range dependencies within the dataset and significantly less biased towards local textures than CNNs [13, 47].
+Consequently, we do not conduct further experiments with other pre-trained weights on the three CNNs.
+Appendix B More quantitative results
+B.1 Detailed Results
+In Tab.
+10, we present the class-wise evaluation results of FLARES-boosted approaches compared to various LiDAR semantic segmentation methods on SemanticKITTI [3], including point-wise, voxel-wise, hybrid, and range-wise approaches.
+In Tab.
+11, we additionally demonstrate the results from nuScenes [16] leaderboard.
+Due to a limited number of entries, we do not categorize methods by publication year and evaluate only a single model using FLARES.
+Clearly, FLARES enables the model to achieve superior accuracy compared to most methods and delivers performance comparable to the best-performing approach, all while maintaining significantly higher computational efficiency.
+Since range-view methods generally offer greater efficiency, it is valuable to compare performance without the use of test-time augmentation.
+We summarize the results in Tab.
+12.
+Rather than comparing with a broad set of range-view approaches introduced over the past decade, we focus on a selection of the most recent and representative methods, as they demonstrate the best performance among them.
+Although FLARES-boosted networks may not exhibit significantly superior segmentation accuracy compared to the latest approaches, they excel in computational efficiency and usage flexibility.
+For instance, RangeFormer [25] requires high computational resources due to its multiple Vision Transformer blocks [13], which contain a substantial number of model parameters.
+Similarly, TFNet [28] processes a sequence of LiDAR data rather than a single frame to incorporate temporal features, and its performance is highly dependent on the sequence length.
+B.2 RangeFormer
+To further asses the potential of our method in approaching the state-of-the-art performance, we reproduced the code for RangeFormer [25], the current best-performing range-view method.
+Results are shown in Tab.
+6.
+Due to the absence of official code, minor discrepancies exist between our reproduced results and the original paper.
+Nonetheless, integrating our method significantly boosts inference speed while achieving comparable accuracy to SOTA methods of other representations.
+Model     Params.     S.KITTI val     nuScenes val     Lat.
+*RangeFormer     24.3M     68.1     77.1     -
+RangeFormer     -     67.0     76.5     87 ms
++FLARES     24.1M     68.9     78.2     55 ms
+Cylinder3D [66]     -     65.9     76.1     -
+SphereFormer [26]     -     67.8     78.4     -
+PVKD [21]     -     66.4     76.0     -
+Table 6: *: Results reported in the original paper [20].
+B.3 Comparison of Data Augmentation
+We compare our two data augmentation methods with existing approaches that share the same objective.
+The results are presented in Tables 8 and 8, evaluated on the val set of the SemanticKITTI dataset [3].
+Aug.     mIoU
+RangePaste [25]     65.4
+Mix3D [35]     64.6
+Ins. CutMix [57]     66.2
+WPD+     67.5
+Table 7: Augmentation methods for class imbalance.
+Aug.     mIoU
+RangeUnion [25]     66.1
+RangeIP [58]     66.9
+MCF     67.5
+Table 8: Augmentation methods for projection noise.
+B.4 Do other standard hyper-parameters matter?
+For consistency, we unify the hyperparameter settings across all experiments.
+To ensure that improvements in segmentation accuracy are not simply due to changes in hyperparameters, we additionally retrain each network using the default settings specified in their respective original papers.
+The corresponding results are reported in Table 9.
+Model     Baseline     FLARES
+    mIoU Lat.     mIoU Lat.
+CENet [5]     62.6 44 ms     65.7 24 ms
+FIDNet [49]     58.9 46 ms     63.4 26 ms
+SalsaNext [7]     59.0 51 ms     64.1 29 ms
+Table 9: We retrain the networks with default parameter settings of the original baselines on SemanticKITTI val set.
+These include batch size, optimizer, learning rate, weight decay, training epochs and loss weights.
+ SemanticKITTI test set
+Method (year)
+mIoU
+car
+bicy
+moto
+truc
+o.veh
+ped
+b.list
+m.list
+road
+park
+walk
+o.gro
+build
+fenc
+veg
+trun
+terr
+pole
+sign
+RandLA-Net [23] [’20]    $50.3$     $94.0$ $19.8$ $21.4$ $42.7$ $38.7$ $47.5$ $48.8$ $4.6$ $90.4$ $56.9$ $67.9$ $15.5$ $81.1$ $49.7$ $78.3$ $60.3$ $59.0$ $44.2$ $38.1$
+PolarNet [63] [’20]    $54.3$     $93.8$ $40.3$ $30.1$ $22.9$ $28.5$ $43.2$ $40.2$ $5.6$ $90.8$ $61.7$ $74.4$ $21.7$ $90.0$ $61.3$ $84.0$ $65.5$ $67.8$ $51.8$ $57.5$
+SqSegV3 [56] [’20]    $55.9$     $92.5$ $38.7$ $36.5$ $29.6$ $33.0$ $45.6$ $46.2$ $\mathbf{20.1}$ $91.7$ $63.4$ $74.8$ $26.4$ $89.0$ $59.4$ $82.0$ $58.7$ $65.4$ $49.6$ $58.9$
+KPConv [46] [’20]    $58.8$     $96.0$ $32.0$ $42.5$ $33.4$ $44.3$ $61.5$ $61.6$ $11.8$ $88.8$ $61.3$ $72.7$ $31.6$ $\mathbf{95.0}$ $64.2$ $84.8$ $69.2$ $69.1$ $56.4$ $47.4$
+FusionNet [40] [’20]    $61.3$     $95.3$ $47.5$ $37.7$ $41.8$ $34.5$ $59.5$ $56.8$ $11.9$ $91.8$ $68.8$ $77.1$ $30.8$ $92.5$ $\mathbf{69.4}$ $\mathbf{85.6}$ $69.8$ $68.5$ $60.4$ $66.5$
+AMVNet [29] [’20]    $\mathbf{65.3}$     $\mathbf{96.2}$ $\mathbf{59.9}$ $54.2$ $48.8$ $45.7$ $\mathbf{71.0}$ $65.7$ $11.0$ $90.1$ $\mathbf{71.0}$ $75.8$ $\mathbf{32.4}$ $92.4$ $69.1$ $\mathbf{85.6}$ $\mathbf{71.7}$ $\mathbf{69.6}$ $\mathbf{62.7}$ $\mathbf{67.2}$
+ΥSalsaNextΥ [10] [’20]    $64.8$     $95.1$ $55.5$ $\mathbf{56.5}$ $\mathbf{60.1}$ $\mathbf{53.7}$ $69.6$ $\mathbf{74.1}$ $11.4$ $\mathbf{93.0}$ $68.9$ $\mathbf{78.9}$ $20.4$ $91.1$ $67.6$ $82.0$ $66.7$ $65.0$ $58.1$ $64.1$
+MPF [1] [’21]    $55.5$     $93.4$ $30.2$ $38.3$ $26.1$ $28.5$ $48.1$ $46.1$ $18.1$ $90.6$ $62.3$ $74.5$ $30.6$ $88.5$ $59.7$ $83.5$ $59.7$ $69.2$ $49.7$ $58.1$
+KPRNet [24] [’21]    $63.1$     $95.5$ $54.1$ $47.9$ $23.6$ $42.6$ $65.9$ $65.0$ $16.5$ $\mathbf{93.2}$ $\mathbf{73.9}$ $\mathbf{80.6}$ $30.2$ $91.7$ $68.4$ $\mathbf{85.7}$ $69.8$ $71.2$ $58.7$ $64.1$
+LiteHDSeg [41] [’21]    $63.8$     $92.3$ $40.0$ $55.4$ $37.7$ $39.6$ $59.2$ $71.6$ $\mathbf{54.3}$ $93.0$ $68.2$ $78.3$ $29.3$ $91.5$ $65.0$ $78.2$ $65.8$ $65.1$ $59.5$ $67.7$
+JS3C-Net [59] [’21]    $66.0$     $95.8$ $59.3$ $52.9$ $54.3$ $46.0$ $69.5$ $65.4$ $39.9$ $88.9$ $61.9$ $72.1$ $31.9$ $\mathbf{92.5}$ $\mathbf{70.8}$ $84.5$ $69.8$ $67.9$ $60.7$ $\mathbf{68.7}$
+Cylinder3D [66] [’21]    $\mathbf{68.9}$     $\mathbf{97.1}$ $\mathbf{67.6}$ $\mathbf{63.8}$ $50.8$ $58.5$ $\mathbf{73.7}$ $69.2$ $48.0$ $92.2$ $65.0$ $77.0$ $32.3$ $90.7$ $66.5$ $85.6$ $\mathbf{72.5}$ $69.8$ $\mathbf{62.4}$ $66.2$
+ΥFIDNetΥ [65] [’21]    $67.4$     $95.8$ $56.7$ $60.7$ $\mathbf{58.1}$ $\mathbf{60.3}$ $72.5$ $\mathbf{72.9}$ $15.8$ $\mathbf{93.2}$ $69.2$ $79.9$ $\mathbf{34.2}$ $91.9$ $69.0$ $84.6$ $68.7$ $\mathbf{70.3}$ $59.9$ $66.9$
+Meta-RSeg [48] [’22]    $61.0$     $93.9$ $50.1$ $43.8$ $43.9$ $43.2$ $63.7$ $53.1$ $18.7$ $90.6$ $64.3$ $74.6$ $29.2$ $91.1$ $64.7$ $82.6$ $65.5$ $65.5$ $56.3$ $64.2$
+PCSCNet [36] [’22]    $62.7$     $95.7$ $48.8$ $46.2$ $36.4$ $40.6$ $55.5$ $68.4$ $\mathbf{55.9}$ $89.1$ $60.2$ $72.4$ $23.7$ $89.3$ $64.3$ $84.2$ $68.2$ $68.1$ $60.5$ $63.9$
+GFNet [39] [’22]    $65.4$     $96.0$ $53.2$ $48.3$ $31.7$ $47.3$ $62.8$ $57.3$ $44.7$ $\mathbf{93.6}$ $\mathbf{72.5}$ $\mathbf{80.8}$ $31.2$ $\mathbf{94.0}$ $\mathbf{73.9}$ $85.2$ $71.1$ $69.3$ $61.8$ $68.0$
+MaskRange [19] [’22]    $66.1$     $94.2$ $56.0$ $55.7$ $59.2$ $52.4$ $67.6$ $64.8$ $31.8$ $91.7$ $70.7$ $77.1$ $29.5$ $90.6$ $65.2$ $84.6$ $68.5$ $69.2$ $60.2$ $66.6$
+GASN [62] [’22]    $\mathbf{70.7}$     $\mathbf{96.9}$ $\mathbf{65.8}$ $58.0$ $\mathbf{59.3}$ $\mathbf{61.0}$ $\mathbf{80.4}$ $\mathbf{82.7}$ $46.3$ $89.8$ $66.2$ $74.6$ $30.1$ $92.3$ $69.6$ $\mathbf{87.3}$ $\mathbf{73.0}$ $\mathbf{72.5}$ $\mathbf{66.1}$ $\mathbf{71.6}$
+ΥCENetΥ [7] [’22]    $68.0$     $95.9$ $61.1$ $\mathbf{62.1}$ $57.2$ $59.0$ $77.2$ $74.2$ $12.2$ $92.2$ $69.9$ $78.7$ $\mathbf{32.9}$ $91.8$ $68.8$ $84.7$ $71.3$ $69.9$ $62.9$ $70.3$
+Table 10: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of SemanticKITTI [3].
+All IoU score are given in percentage (%).
+All approaches are categorized by the year of publication.
+In each block, bold and underline indicate the best and second best result in the column.
+ nuScenes test set
+Method (year)
+mIoU
+barrier
+bicy
+bus
+car
+const
+moto
+ped
+traffic.c
+trailer
+truck
+driv
+o.flat
+side
+terrain
+manm
+veg
+PolarNet [63] [’20]    $69.4$     $72.2$ $16.8$ $77.0$ $86.5$ $51.1$ $69.7$ $64.8$ $54.1$ $69.7$ $63.5$ $96.6$ $67.1$ $77.7$ $72.1$ $87.1$ $84.5$
+JS3C-Net [59] [’21]    $73.6$     $80.1$ $26.2$ $87.8$ $84.5$ $55.2$ $72.6$ $71.3$ $66.3$ $76.8$ $71.2$ $96.8$ $64.5$ $76.9$ $74.1$ $87.5$ $86.1$
+AMVNet [29] [’20]    $77.3$     $80.6$ $32.0$ $81.7$ $88.9$ $67.1$ $84.3$ $76.1$ $73.5$ $\mathbf{84.9}$ $67.3$ $97.5$ $67.4$ $79.4$ $75.5$ $\mathbf{91.5}$ $88.7$
+Cylinder3D [66] [’21]    $77.2$     $82.8$ $29.8$ $84.3$ $89.4$ $63.0$ $79.3$ $\mathbf{77.2}$ $73.4$ $84.6$ $69.1$ $\mathbf{97.7}$ $70.2$ $\mathbf{80.3}$ $75.5$ $90.4$ $87.6$
+RangeFormer [25] [’23]    $\mathbf{80.1}$     $\mathbf{85.6}$ $\mathbf{47.4}$ $\mathbf{91.2}$ $\mathbf{90.9}$ $\mathbf{70.7}$ $\mathbf{84.7}$ $77.1$ $\mathbf{74.1}$ $83.2$ $\mathbf{72.6}$ $97.5$ $\mathbf{70.7}$ $79.2$ $75.4$ $91.3$ $88.9$
+ΥCENetΥ [’22]    $77.5$     $82.2$ $31.1$ $85.2$ $90.7$ $65.1$ $82.5$ $74.9$ $72.3$ $84.7$ $70.5$ $97.4$ $69.5$ $79.6$ $\mathbf{76.1}$ $90.5$ $\mathbf{90.1}$
+Table 11: The class-wise IoU scores compared with different LiDAR semantic segmentation approaches on the test set of nuScenes [16].
+All IoU score are given in percentage (%).
+Due to limited number of submission times, we only test FLARES with CENet [7].
+Bold and underline indicate the best and second best result in the column.
+ SemanticKITTI test set
+Method (year)     #params
+mIoU
+car
+bicy
+moto
+truc
+o.veh
+ped
+b.list
+m.list
+road
+park
+walk
+o.gro
+build
+fenc
+veg
+trun
+terr
+pole
+sign
+RangeFormer [25] [’23]     24.3M     $\mathbf{69.5}$     $94.7$ $60.0$ $\mathbf{69.7}$ $\mathbf{57.9}$ $\mathbf{64.1}$ $72.3$ $\mathbf{72.5}$ $\mathbf{54.9}$ $90.3$ $69.9$ $74.9$ $\mathbf{38.9}$ $90.2$ $66.1$ $\mathbf{84.1}$ $68.1$ $\mathbf{70.0}$ $58.9$ $63.1$
+LENet [12] [’23]     4.7M     $64.5$     $93.9$ $57.0$ $51.3$ $44.3$ $44.4$ $66.6$ $64.9$ $36.0$ $91.8$ $68.3$ $76.9$ $30.5$ $91.2$ $66.0$ $83.7$ $68.3$ $67.8$ $58.6$ $63.2$
+TFNet [28] [’24]     -     $66.1$     $94.3$ $\mathbf{60.7}$ $58.5$ $38.4$ $48.4$ $74.3$ $72.2$ $35.5$ $90.6$ $68.5$ $75.3$ $29.0$ $\mathbf{91.6}$ $67.3$ $83.8$ $\mathbf{71.1}$ $67.0$ $\mathbf{60.8}$ $\mathbf{68.7}$
+ΥSalsaNextΥ [’20]     6.7M     $63.3$     $94.7$ $52.9$ $55.7$ $57.3$ $50.2$ $65.5$ $70.9$ $13.0$ $\mathbf{92.6}$ $69.0$ $77.7$ $20.5$ $90.4$ $65.8$ $80.8$ $65.0$ $63.4$ $55.4$ $62.4$
+ΥFIDNetΥ [’21]     6.1M     $65.1$     $95.3$ $51.0$ $57.0$ $54.8$ $58.1$ $68.1$ $68.9$ $14.4$ $92.3$ $68.3$ $\mathbf{78.0}$ $32.3$ $\mathbf{91.6}$ $\mathbf{67.6}$ $83.7$ $66.6$ $68.8$ $55.1$ $64.8$
+ΥCENetΥ [’22]     6.8M     $66.6$     $\mathbf{95.6}$ $58.5$ $61.6$ $51.7$ $50.2$ $\mathbf{74.5}$ $72.4$ $23.2$ $91.4$ $69.6$ $77.1$ $31.7$ $91.1$ $66.6$ $83.8$ $69.9$ $68.3$ $60.3$ $\mathbf{68.7}$
+ΥRangeViTΥ [’23]     23.6M     $66.1$     $\mathbf{95.6}$ $56.3$ $60.5$ $52.4$ $\mathbf{57.1}$ $72.0$ $69.7$ $16.0$ $91.6$ $\mathbf{71.1}$ 77.3{77.3} $32.7$ $91.4$ $67.4$ $83.1$ $68.0$ $68.1$ $58.0$ $67.5$
+Table 12: The class-wise IoU scores compared with latest range-view LiDAR semantic segmentation approaches on the test set of SemanticKITTI [3].
+All IoU score are given in percentage (%).
+All approaches are evaluated without test-time augmentation.
+Bold and underline indicate the best and second best result in the column.
+Appendix C More qualitative results
+[…]
+Figure 14: Additional qualitative results of SalsaNext [10]: points in red and gray represent incorrect and correct predictions, respectively.
+The groundtruth image is color-coded to differentiate between classes.
+Figure 15: Additional qualitative results of FIDNet [65]: points in red and gray represent incorrect and correct predictions, respectively.
+The groundtruth image is color-coded to differentiate between classes.
+Figure 16: Additional qualitative results of CENet [7]: points in red and gray represent incorrect and correct predictions, respectively.
+The groundtruth image is color-coded to differentiate between classes.
+Figure 17: Additional qualitative results of RangeViT [2]: points in red and gray represent incorrect and correct predictions, respectively.
+The groundtruth image is color-coded to differentiate between classes.
+Appendix D Video Demo
+In addition to the figures, we have included two video demos per network tested (eight videos in total) in the supplementary materials to showcase a more comprehensive evaluation of our approach.
+Each video contains hundreds of frames, with each frame presenting three visualizations: groundtruth, baseline prediction, and FLARES prediction.
+Each visualization includes both top-down-view and range-view images of the LiDAR point cloud.
+For videos labeled with the suffix errormap, predictions are displayed in a binary format, with incorrect predictions shown in red and correct ones in gray.